### v1

# Quantifying the effectiveness of linear preconditioning in Markov chain Monte Carlo

Abstract
Linear transformation of the state variable (linear preconditioning) is a common technique that often drastically improves the practical performance of a Markov chain Monte Carlo algorithm. Despite this, however, the benefits of linear preconditioning are not well-studied theoretically, and rigorous guidelines for choosing preconditioners are not always readily available. Mixing time bounds for various samplers have been produced in recent works for the class of strongly log-concave and Lipschitz target distributions and depend strongly on a quantity known as the condition number. We study linear preconditioning for this class of distributions, and under appropriate assumptions we provide bounds on the condition number after using a given linear preconditioner. We provide bounds on the spectral gap of RWM that are tight in their dependence on the condition number under the same assumptions. Finally we offer a review and analysis of popular preconditioners. Of particular note, we identify a surprising case in which preconditioning with the diagonal of the target covariance can actually make the condition number increase relative to doing no preconditioning at all.
Keywords Markov chain Monte Carlo  $\cdot$ Preconditioning  $\cdot$ Bayesian inference  $\cdot$ Bayesian Computation  $\cdot$ Condition Number
1 Introduction
Markov chain Monte Carlo (MCMC) techniques are ubiquitous in the field of Bayesian computation. Decades of use have proven their effectiveness and longevity in the face of major changes to the structure of statistical models. To keep pace with the increasingly complex problems faced by practitioners, the MCMC community has seen the invention of variants of the original Metropolis-Hastings samplers (Metropolis et al., 1953; Hastings, 1970) such as the Metropolis adjusted Langevin algorithm (MALA) (Roberts and Tweedie, 1996; Roberts and Rosenthal, 1998) and Hamiltonian Monte Carlo (HMC) (Neal, 2011).
Preconditioning was conceived as a way to improve the stability of iterative solvers in linear algebra. The conditioning of a particular problem corresponds to how easily it can be solved, and is quantified by a condition number. The term ‘condition number’ was first coined by Turing (1948), although similar quantities are used in earlier works (Wittmeyer, 1936; von Neumann and Goldstine, 1947). In its most basic form one can think of preconditioning as transforming to a space in which the problem at hand is both easier to solve and more amenable to implementation on a computer. Assuming that the transformation is invertible, the idea is to solve the problem in the transformed space and then transform the solution back to the original space using the inverse.
Recent works by Dalalyan (2017); Durmus and Moulines (2017); Chen et al. (2020); Mangoubi and Smith (2021); Chewi et al. (2021); Lee et al. (2021); Andrieu et al. (2022) and others establish non-asymptotic mixing time bounds for various MCMC samplers. The choice of distance metric $\mathcal{D}$ can vary, but excluding polylogarithmic terms each bound can be written in the form
$\inf\{n>0:\mathcal{D}(P^{n},\Pi)<\epsilon\}\leq C_{\epsilon}\kappa^{\alpha}d^{\beta},$
where $P^{n}$ denotes the marginal distribution of the $n$th state, $\Pi$ is the equilibrium distribution, $d$ is the dimension of the state space, $\alpha$ and $\beta$ are positive constants and $C_{\epsilon}>0$ is an explicit constant depending on $\epsilon>0$ (see Table 1 for more detail). The quantity $\kappa$ (defined in equation (4)) has come to be known as the condition number within the MCMC community. Lee et al. (2021) show that for MALA and HMC there are strongly log-concave distributions for which $\alpha\geq 1$, and the best known bounds to our knowledge for these methods set $\alpha\geq 1$.
In this paper we study how the common practice of preconditioning an MCMC sampler affects the condition number $\kappa$. When the transformation is linear we refer to it as linear preconditioning. Although it is widely used in practice, the effects of linear preconditioning in MCMC are largely under-explored theoretically (though see Roberts and Rosenthal (2001) for some asymptotic results in particular cases).
1.1 Notation
We use $\mathcal{X}$ to denote the Borel $\sigma$-field of $\mathbb{R}^{d}$. The set of non-negative real numbers is denoted by $\mathbb{R}^{+}$. Let $\Pi$ be a measure on $(\mathbb{R}^{d},\mathcal{X})$ with density $\pi:\mathbb{R}^{d}\to\mathbb{R}^{+}$ with respect to the Lebesgue measure on $\mathbb{R}^{d}$. The potential associated with the density $\pi$ is the function $U:\mathbb{R}^{d}\to\mathbb{R}$ satisfying the relation $\pi(x)\propto\exp(-U(x))$. We will often refer to the measure that MCMC algorithms are intended to sample from as the target. The normal distribution with mean $\mu\in\mathbb{R}^{d}$ and covariance $\Sigma\in\mathbb{R}^{d\times d}$ is $\mathcal{N}(\mu,\Sigma)$ and its density at $x\in\mathbb{R}^{d}$ is $\mathcal{N}(x;\mu,\Sigma)$.
We denote by $L^{2}(\Pi)$ the Hilbert space of functions that are square integrable with respect to $\Pi$ with inner product
[…]
We denote by $L_{0}^{2}(\Pi)$ the subset of $L^{2}(\Pi)$ containing the functions with expectation 0 with respect to $\Pi$. When $P$ is a $\Pi$-invariant Markov Kernel we define an associated bounded linear operator $P:L^{2}(\Pi)\to L^{2}(\Pi)$ as $Pf(x):=\int_{\mathbb{R}^{d}}P(x,dy)f(y)$. We denote by $I$ the identity mapping on $L^{2}(\Pi)$. For a given operator $P$ we define the Dirichlet form $\mathcal{E}(P,f):=\langle(I-P)f,f\rangle$ for all $f\in L^{2}(\Pi)$. When $P$ is $\Pi$-reversible we define its right spectral gap (which we will from this point refer to as its spectral gap) as
$\gamma:=\inf_{f\in L_{0}^{2}(\Pi),f\not\equiv 0}\frac{\mathcal{E}(P,f)}{\text{Var}_{\Pi}(f)}$
The set $\{1,2,...,k\}$ is denoted as $[k]$ for any positive integer $k$. For a function $g\in C^{2}(\mathbb{R}^{d})$ we define $\nabla g(y)\in\mathbb{R}^{d}$ and $\nabla^{2}g(y)\in\mathbb{R}^{d\times d}$ elementwise as
$\left(\nabla g(y)\right)_{i}:=\frac{\partial}{\partial y_{i}}g(y)\text{ }\left(\nabla^{2}g(y)\right)_{ij}:=\frac{\partial^{2}}{\partial y_{i}\partial y_{j}}g(y)$
for $i,j\in[d]$.
[…]
$\kappa(A):=\frac{\max_{i\in[d]}|\lambda_{i}(A)|}{\min_{i\in[d]}|\lambda_{i}(A)|}$
We overload the diag function as follows: $\text{diag}(A)$ is the diagonal matrix that shares its diagonal with $A\in\mathbb{R}^{d\times d}$ and $\text{diag}\{f(i):i\in[d]\}\in\mathbb{R}^{d\times d}$ is the diagonal matrix whose $(i,i)$th element is $f(i)$. We denote by $GL_{d}(\mathbb{R})$ the set of invertible $d\times d$ matrices over $\mathbb{R}$. We use the $\preceq$’ and $\succeq$’ relations to take symmetric matrices in the following way: $A\preceq B$ (resp. $A\succeq B$) if and only if $B-A$ (resp. $A-B$) is positive semidefinite. The ordering generated by these relations is known as the Loewner order. The relations $\prec$’ and $\succ$’ are defined similarly, replacing the semidefiniteness condition with definiteness.
For two real valued functions $f(n)$ and $g(n)$ we say $f(n)=O(g(n))$ if there exists a universal constant $K>0$ such that $f(n)\leq Kg(n)$. We say $f(n)=\Omega(g(n))$ if there exists a universal constant $K>0$ such that $f(n)\geq Kg(n)$.
1.2 Main contributions
We begin by showing that not all distributions are amenable to linear preconditioning, even within the class of strongly log-concave distributions (see Proposition 1). We then establish upper bounds on the condition number after preconditioning with $L\in GL_{d}(\mathbb{R})$ under certain conditions on $\nabla^{2}U$ and $L$ in Sections 3.2 and 3.3. In Section 3.2 we initially we impose conditions on the eigenvalues and eigenvectors of $\nabla^{2}U$ and $L$, but then show how these can be controlled using only knowledge of the norm $\|\nabla^{2}U-LL^{T}\|$ and an ‘eigengap’ condition, using techniques from matrix perturbation theory. We then further relax the eigengap condition, providing bounds when only the norm can be controlled. See Theorems 3, 4, and 5 for a thorough treatment. We demonstrate that our results naturally apply to models $\Pi$ whose Hessians $\nabla^{2}U$ can be written in the form $A+B(x)$ such that $\|B(x)\|$ is small, and refer to this as the additive Hessian structure, see (8). In Section 3.3 we establish upper bounds on the condition number after linearly preconditioning models whose Hessians have a multiplicative structure (defined in equation (9)) as is the case for generalised linear models. Our results can be directly inserted into various mixing time bounds that have been established for MCMC samplers such as MALA and HMC to see how much the mixing time can be reduced after linear preconditioning.
Andrieu et al. (2022) establish both upper and lower bounds on the spectral gap of the random walk Metropolis that have the same order dependence on dimension. The lower bound is explicit in the condition number, but the upper bound is not. Our next contribution is therefore to make the upper bound explicit in the condition number under verifiable conditions on $\nabla^{2}U$. The resulting bounds are both explicit in, and matched in their dependence on the condition number. Specifically, where $\gamma_{\kappa}$ is the spectral gap of the random walk Metropolis on a model with condition number $\kappa$ the bounds are
$C\xi\exp(-2\xi)\cdot\frac{1}{\kappa}\cdot\frac{1}{d}\leq\gamma_{\kappa}\leq(1+2\epsilon)\frac{\xi}{2}\cdot\frac{1}{\kappa}\cdot\frac{1}{d}$
where $C=1.972\times{10}^{-4}$, $\xi>0$ is a constant depending on the step-size, and $\epsilon>0$ is a constant. See Theorem 11 for more details. Combining this with our condition number upper bounds, we are able to provide concrete scenarios in which this spectral gap of the random walk Metropolis provably increases under linear preconditioning. See Corollary 12 for particulars.
Finally we evaluate various possible choices of linear preconditioner $L$. We identify a surprising case in which diagonal preconditioning with $L=\text{diag}(\Sigma_{\pi})^{-1/2}\in\mathbb{R}^{d\times d}$ (where $\Sigma_{\pi}$ is the target covariance) increases the condition number instead of decreasing it. See Section 3.5.3. The result is particularly relevant as diagonal preconditioning in this way is the default approach in many popular MCMC software packages. We then consider choices such as $L=\Sigma_{\pi}^{-1/2}$ and $L=(\nabla^{2}U(x^{*}))^{1/2}$, where $x^{*}$ is the mode of $\Pi$, as well as others proposed by Dalalyan (2017) and Titsias (2023), and in each case offer conditions on $\nabla^{2}U$ under which our theoretical results can be applied. We also discuss situations in which certain choices might be preferred to others.
2 Preconditioning in MCMC
Suppose $\Pi$ is a measure on $(\mathbb{R}^{d},\mathcal{X})$ with density $\pi(x)\propto e^{-U(x)}$ for some potential $U:\mathbb{R}^{d}\to\mathbb{R}$. The aim of MCMC is to sample from $\Pi$ so as to infer its properties. In particular we are interested in estimating expectations $\mathbb{E}_{\pi}[f(X)]$ of some functional $f$. Given samples $\{X_{i}\}_{i=1}^{N}$ we achieve this by forming the estimator $\hat{f}_{N}:=N^{-1}\sum_{i=1}^{N}f(X_{i})$. In many situations of interest drawing independent samples from $\Pi$ is computationally infeasible. The lifeblood of MCMC is in creating and analyzing algorithms which produce Markov chains whose equilibrium distribution is $\Pi$. We can then use the samples from these Markov chains in estimators such as $\hat{f}_{N}$.
2.1 MCMC
Constructing a Markov chain with an arbitrary equilibrium distribution is seemingly a highly nontrivial task. However the Metropolis-Hastings (M-H) filter (Metropolis et al., 1953; Hastings, 1970) allows us to do just that. Given a Markov kernel $Q_{\theta}(x\to.):\mathcal{X}\to\mathbb{R}^{+}$ and its density $q_{\theta}(x\to.)$ we can construct a $\Pi$-invariant Markov chain using the pseudocode in Algorithm 1. To apply step 2 to an existing Markov kernel $Q_{\theta}$ is to Metropolize it. Most of the algorithms considered in this paper appear in their Metropolized form.
input : Chain length $N$, Initial distribution $\mu$, Initial state $X_{0}\sim\mu$, Proposal parameters $\theta$
output : $\Pi$-invariant Markov chain $\{X_{i}\}_{i=1}^{N}$
for $i\in\{0,...,N-1\}$ do
  1. 1.
    Proposal: Propose a new state $X^{\prime}_{i}\sim Q_{\theta}(X_{i}\to.)$.
  2. 2.
    M-H filter: Draw $U\sim U[0,1]$ and calculate
    $\alpha(X_{i}\to X^{\prime}_{i}):=\min\left\{1,\frac{\pi(X^{\prime}_{i})q_{\theta}(X^{\prime}_{i}\to X_{i})}{\pi(X_{i})q_{\theta}(X_{i}\to X^{\prime}_{i})}\right\}$
    If $U\leq\alpha(X_{i}\to X^{\prime}_{i})$ set $X_{i+1}=X^{\prime}_{i}$, otherwise set $X_{i+1}=X_{i}$.
end for
Algorithm 1 Metropolized Markov chain
When $q_{\theta}(x\to.):=\mathcal{N}(\hskip 2.84544pt.\hskip 2.84544pt;x,\sigma^{2}\mathbf{I}_{d})$ Algorithm 1 defines the random walk Metropolis (RWM) algorithm. Here $\sigma>0$ is a step-size and must be tuned by the practitioner. When $q_{\theta}(x\to.):=\mathcal{N}(\hskip 2.84544pt.\hskip 2.84544pt;x-\sigma^{2}\nabla U(x)/2,\sigma^{2}\mathbf{I}_{d})$ Algorithm 1 defines the Metropolis-adjusted Langevin algorithm (MALA). Note that this proposal mechanism follows a single step with step-size $\sigma^{2}$ of the Euler–Maruyama discretisation of the overdamped Langevin diffusion, described through the dynamics
$dX_{t}=-\frac{1}{2}\nabla U(X_{t})dt+dB_{t},$ (1)
where $B_{t}$ is a standard Brownian motion on $\mathbb{R}^{d}$ (we drop the word overdamped from this point forward). The Langevin diffusion associated with a given potential $U$ is an object of independent interest since it admits $\Pi$ as equilibrium distribution under mild conditions (e.g. Roberts and Tweedie (1996)).
The idea underpinning HMC (Duane et al., 1987; Neal, 1996) is to extend the state space to $\mathbb{R}^{d}\times\mathbb{R}^{d}$ to include a momentum vector, viewing the states in the Markov chain as the positions in physical space of a particle moving in the potential $U$. To propose a new state $X^{\prime}_{i}$ we simply evolve a particle with initial position $X_{i}$ according to a discretisation of Hamiltonian dynamics with random initial momentum and potential $U$. Specifically we propose $X^{\prime}_{i}=x_{T}$ using the final value $(x_{T},p_{T})$ of the trajectory described by the Hamiltonian system:
$\frac{dx_{t}}{dt}=\nabla_{p}H(x_{t},p_{t})\text{, }\frac{dp_{t}}{dt}=-\nabla_{x}H(x_{t},p_{t})$ (2)
with initial conditions $x_{0}=X_{i}$, $p_{0}\sim\mathcal{N}(0,\mathbf{I}_{d})$. The function $H:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ is the Hamiltonian $H(x,p):=U(x)+K(p)$ where $K(p)=p^{T}p/2$ is the kinetic energy. Note that if we marginalise the density proportional to $\exp(-H(x,p))$ over $p$ we are left with $\pi$. See Neal (2011); Betancourt et al. (2017) for a more thorough exposition, Livingstone et al. (2019a) for extensions to different forms of $K(p)$, and Chen et al. (2014) for a stochastic gradient version of the algorithm. Convergence properties are discussed in Durmus et al. (2020); Livingstone et al. (2019b). For a more general discussion of schemes involving augmentation of the state vector see Andrieu et al. (2020); Glatt-Holtz et al. (2020).
2.2 Well-conditioned distributions and the condition number
2.2.1 Background assumptions on $\Pi$
Here we introduce a background assumption which is assumed for most theorems in the text.
Assumption 2.1.
The potential $U\in C^{2}$, and there exist constants $M\geq m>0$ such that
$m\mathbf{I}_{d}\preceq\nabla^{2}U(x)\preceq M\mathbf{I}_{d}$ (3)
for all $x\in\mathbb{R}^{d}$.
The lower bound on the Hessian in Assumption 2.1 is known as $m$-strong convexity of $U$ and the upper bound is commonly known as $M$-smoothness of $U$. Note that $M$-smoothness holds if and only if $\nabla U$ is $M$-Lipschitz.
2.2.2 The condition number
When $U$ satisfies Assumption 2.1 the condition number is defined as
$\kappa:=\frac{M}{m}.$ (4)
When $\Pi$ is Gaussian with covariance $\Sigma_{\pi}\in\mathbb{R}^{d\times d}$, $\kappa=\lambda_{1}(\Sigma_{\pi})/\lambda_{d}(\Sigma_{\pi})$, which is the spectral condition number of $\Sigma_{\pi}$. When $U\in C^{2}$ the condition number can be equivalently defined as
$\kappa:=\sup_{x\in\mathbb{R}^{d}}\|\nabla^{2}U(x)\|\sup_{x\in\mathbb{R}^{d}}\|\nabla^{2}U(x)^{-1}\|$ (5)
It is bounded below by 1, and a larger condition number traditionally signifies a harder problem that is less amenable to implementation on a computer. The significance of $\kappa$ in the context of sampling is demonstrated by its presence in bounds on quantities which govern the performance of MCMC algorithms such as the relaxation time and the $\epsilon$-mixing time. We present a selection of these bounds in Table 1. Each is polynomial in both the dimension and the condition number. The selection is not exhaustive, we merely present it to highlight the ubiquity of the condition number.
Relaxation Time $\epsilon$-Mixing Time
$\tilde{O}\left(\kappa^{\frac{4}{3}}d^{\frac{1}{2}}\log\frac{1}{\epsilon}\right)^{\bigtriangledown}$ Chewi et al. (2021)
Upper Bound $O(\kappa d)^{\bigtriangleup}$ Andrieu et al. (2022) $\tilde{O}\left(\kappa d\log\frac{1}{\epsilon}\right)^{\bigtriangleup}$ Andrieu et al. (2022)
$\tilde{O}\left(\kappa d^{\frac{2}{3}}\log\frac{1}{\epsilon}\right)^{\bigcirc}$ Chen et al. (2020)
Lower Bound $\Omega\left(\frac{\kappa d}{\log d}\right)^{\bigtriangledown}$ Lee et al. (2021) $\Omega\left(\frac{\kappa d}{\log^{2}d}\right)^{\bigtriangledown}$ Lee et al. (2021)
[…]
There also exist numerous bounds on the relaxation and $\epsilon$-mixing time of the unmetropolized counterparts of the algorithms in Table 1 (for example, see Mangoubi and Vishnoi (2018); Dalalyan (2017).
2.3 Preconditioning
Assuming that $X\sim\Pi$ we let $\tilde{E}:=(\mathbb{R}^{d},\mathcal{X})$ be the space generated by the diffeomorphism $g:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $Y:=g(X)$ is distributed according to the pushforward $\tilde{\Pi}:=g_{\#}\Pi$ which is a distribution on $\tilde{E}$ with density $\tilde{\pi}$. Preconditioning is the act of identifying a transformation $g$ such that $\tilde{\Pi}$ admits a faster mixing MCMC algorithm than $\Pi$. Therefore we would like the condition number of $\tilde{\Pi}$ to be significantly lower than that of $\Pi$. We would also like that for two samples $Y$ and $X$ sampled approximately from $\tilde{\Pi}$ and $\Pi$, the increase in sample quality from $X$ to $g^{-1}(Y)$ is not outweighed by the computational complexity of sampling $Y$ and calculating $g^{-1}(Y)$.
3 Linear preconditioning
Linear preconditioning is the act of performing a linear transformation to the state variable in the hopes of reducing the condition number. Concretely, we are seeking the preconditioner $L\in GL_{d}(\mathbb{R})$ which constitutes the transform $Y=LX$, such that the condition number of the distribution of $Y$, which we will call $\kappa_{L}$, is small. From (5) we see that
$\kappa_{L}=\sup_{y\in\mathbb{R}^{d}}\left\|\nabla^{2}\tilde{U}(y)\right\|\sup_{y\in\mathbb{R}^{d}}\left\|\nabla^{2}\tilde{U}(y)^{-1}\right\|=\sup_{x\in\mathbb{R}^{d}}\left\|L^{-T}\nabla^{2}U(x)L^{-1}\right\|\sup_{x\in\mathbb{R}^{d}}\left\|L\nabla^{2}U(x)^{-1}L^{T}\right\|$ (6)
where $\tilde{U}:\mathbb{R}^{d}\to\mathbb{R}$ is the potential associated with $\tilde{\Pi}$. For a concrete example, Figure 1 shows two contour plots. The red contour plot is generated by the posterior of a Bayesian logistic regression, the green plot shows the distribution after preconditioning with an appropriate linear preconditioner.
Figure 1: A red contour plot (left) showing the contours of a Bayesian logistic regression posterior, and a green contour plot (right) showing the contours of the posterior preconditioned with an appropriate matrix $L\in\mathbb{R}^{d\times d}$
Often linear preconditioning is encountered as a modification to the proposal distribution of a canonical MCMC algorithm and not as a transformation. In every case these modifications can be shown to be equivalent to making a linear transformation of the canonical algorithm without preconditioning. For instance, preconditioned MALA has proposal
$Y^{\prime}:=Y+\sigma^{2}A\nabla_{y}\log\tilde{\pi}(Y)+\sqrt{2\sigma^{2}}A^{\frac{1}{2}}\xi$
where $\sigma^{2}>0$ is the step-size, $\tilde{\pi}$ is the target density, $\xi\sim\mathcal{N}(0,\mathbf{I}_{d})$, and $A$ is a positive definite ‘preconditioning matrix’. This is equivalent to making a proposal
$X^{\prime}=X+\sigma^{2}\nabla_{x}\log\pi(X)+\sqrt{2\sigma^{2}}\xi$
under the transformation $Y=LX$ where $L=A^{1/2}$ and $\pi(X)=\tilde{\pi}(Y)|\det A^{1/2}|$. That the acceptance probabilities are equal dictates that the chains $\{Y_{i}\}$ and $\{X_{i}\}$ are isomorphic in the sense of Johnson and Geyer (2012, Appendix A) which means, amongst other things, that they have the same spectral gap. We frame preconditioning in terms of a transformation because it suits our analysis.
3.1 Unpreconditionable Distributions
Methods to adaptively seek such a transformation are implemented in the MCMC samplers provided by the major software packages. For instance the HMC sampler in the popular statistical modelling platform Stan (Stan Development Team, 2023) offers the ability to infer the target covariance $\Sigma_{\pi}\in\mathbb{R}^{d\times d}$ giving an estimator $\hat{\Sigma}_{\pi}\in\mathbb{R}^{d\times d}$ to use as the inverse metric. This is equivalent to linear preconditioning with $L=\hat{\Sigma}_{\pi}^{-1/2}$. Since computational effort is required to infer $\hat{\Sigma}_{\pi}$ the idea is that preconditioning with $L=\hat{\Sigma}_{\pi}^{-1/2}$ is better than doing nothing. As we show in the below proposition, however, even within the class of models satisfying Assumption 2.1, this will not always be the case.
Proposition 1.
There exist distributions satisfying Assumption 2.1 for which any non-orthogonal linear preconditioner will cause the condition number to increase.
One such distribution has density of the form $\pi(x,y)\propto\exp(-U(x,y))$ where
$U(x,y)=\frac{m-M}{2}\left(\cos x+\cos y\right)+\frac{M+m}{2}\left(\frac{x^{2}}{2}+\frac{y^{2}}{2}\right)$ (7)
for $x,y\in\mathbb{R}$. For such a target the condition number $\kappa_{L}$ is bounded below by $\kappa(LL^{T})\kappa$, where $\kappa(LL^{T})$ is the spectral condition number of $LL^{T}$. The Hessian of $U$ is in the form $\text{diag}\{f(x),f(y)\}$ where $f$ ranges freely in $[m,M]$. The lower bound $\kappa(LL^{T})\kappa$ highlights that any preconditioning causes the target to be more ill-conditioned by an amount exactly proportional to $\kappa(LL^{T})$. The full proof of Proposition 1 can be found in Appendix A.1.
The issue for the potential (7) is that every eigenvalue of $\nabla^{2}U$ can assume both the value $m$ and the value $M$ at given points in $\mathbb{R}^{d}$. In the following two sections we characterise effective linear preconditioners for two broad classes of models. We do this by establishing upper bounds on the condition number after linear preconditioning under model appropriate assumptions.
Throughout Section 3.2 and 3.3 $\lambda_{i}(x)$ denotes the $i$-th largest eigenvalue of $\nabla^{2}U(x)$ and $v_{i}(x)$ the corresponding normalised eigenvector. Since the Hessian is everywhere symmetric its eigenvectors are orthogonal for a fixed $x\in\mathbb{R}^{d}$. At some points in this section (and elsewhere in the text) it is useful to assume that $L$ is symmetric. The following proposition shows that this can be done without loss of generality.
Proposition 2.
The condition number after linear preconditioning with a matrix $L\in GL_{d}(\mathbb{R})$ is equivalent to the condition number after linear preconditioning with $\tilde{L}\in GL_{d}(\mathbb{R})$, where $\tilde{L}$ is symmetric, positive definite, and has eigenvalues equal to the singular values of $L$.
A proof can be found in Appendix A.2. If $L$ is assumed to be symmetric we can then denote its $i$-th eigenvalue as $\sigma_{i}$, with associated eigenvector $v_{i}$.
3.2 Linear preconditioning for additive Hessians
We call a Hessian additive if it has the form
$\nabla^{2}U(x)=A+B(x)$ (8)
where $A,B(x)\in\mathbb{R}^{d\times d}$ are symmetric and $\|B(x)\|$ is ‘small’ for all $x\in\mathbb{R}^{d}$. Examples of models whose potentials have Hessians in this form include: Gaussians ($B(x)\equiv 0$), strongly log-concave mixture of Gaussians (Dalalyan, 2017, section 6.1), and Bayesian Huberized regressions with strongly log-concave smooth priors (Rosset and Zhu, 2004). The results in this section are presented in generality, but the assumptions under which they hold are particularly appropriate for models with an additive Hessian.
We first present a general result under the following assumptions on the eigenstructure of $\nabla^{2}U$ and $L$.
Assumption 3.1.
There exists an $\epsilon>0$ such that
$(1+\epsilon)^{-1}\leq\frac{\lambda_{i}(x)}{\sigma_{i}^{2}}\leq 1+\epsilon$
for all $i\in[d]$ and $x\in\mathbb{R}^{d}$.
Assumption 3.2.
There exists a $\delta>0$ such that $v_{i}(x)^{T}v_{i}\geq 1-(1-\sqrt{1-\delta})^{2}$ for all $i\in[d]$ and $x\in\mathbb{R}^{d}$.
Theorem 3.
Let $\Pi$ have potential $U$ satisfying Assumption 2.1. For a given preconditioner $L\in GL_{d}(\mathbb{R})$ for which Assumptions 3.1 and 3.2 hold, the condition number after preconditioning satisfies
$\kappa_{L}\leq(1+\epsilon)^{2}\left(1+\delta\sqrt{\sum_{i=1}^{d}\sigma_{i}^{2}\sum_{i=1}^{d}\sigma_{i}^{-2}}\right)^{4}$
Proof can be found in Appendix A.4. Assumption 3.1 states that the eigenvalues of $\nabla^{2}U(x)$ do not change much over $\mathbb{R}^{d}$. Assumption 3.2 implies that $v_{i}(x)^{T}v_{i}\geq 1-\delta$ for $i\in[d]$, $x,y\in\mathbb{R}^{d}$ and $v_{i}(x)^{T}v_{j}\leq\delta$ for $i\in[d]$, $x,y\in\mathbb{R}^{d}$ where $i\neq j$ (see Appendix A.3 for details). In the Gaussian case $\epsilon=\delta=0$ when $L=\Sigma_{\pi}^{-1/2}$ the bound becomes $\kappa_{L}\leq 1$, as expected.
Remark.
In the case of the additive Hessian, the eigenvalue stability inequality implies that $\lambda_{i}(A)-\|B(x)\|\leq\lambda_{i}(A+B(x))\leq\lambda_{i}(A)+\|B(x)\|$ for all $i\in[d]$. Therefore choosing $L=A^{1/2}$ gives us that
$1-\frac{\|B(x)\|}{\lambda_{d}(A)}\leq\frac{\lambda_{i}(x)}{\sigma_{i}^{2}}\leq 1+\frac{\|B(x)\|}{\lambda_{d}(A)}$
for all $i\in[d]$. So if $\|B(x)\|$ is small, $L=A^{1/2}$ will yield a smaller $\epsilon$. If $\|B(x)\|$ is large but $B(x)$ does not exhibit large variations we can simply restate the Hessian as $\nabla^{2}U(x)=\tilde{A}+\tilde{B}(x)$ where $\tilde{A}:=A+B(x^{*})$ and $\tilde{B}(x):=B(x)-B(x^{*})$ where $x^{*}\in\mathbb{R}^{d}$ is some point in the state space, and use $L=\tilde{A}^{1/2}$.
Remark.
Note that the quantity $\sqrt{\sum_{i}\sigma_{i}^{2}\sum_{i}\sigma_{i}^{-2}}=\sqrt{\text{Tr}(LL^{T})\text{Tr}((LL^{T})^{-1})}$ is upper bounded by $d\sqrt{\kappa(LL^{T})}$. It can be viewed as an alternative ‘condition number’ of $LL^{T}$, similar to that proposed in Langmore et al. (2020).
The fact that $\sqrt{\sum_{i}\sigma_{i}^{2}\sum_{i}\sigma_{i}^{-2}}$ multiplies $\delta$ shows that when sampling from highly anisotropic distributions the penalty for misaligned eigenvectors of $LL^{T}$ relative to $\nabla^{2}U$ is larger. As an example take $\Pi=\mathcal{N}(0,\Sigma_{\pi})$ where $\Sigma_{\pi}\in\mathbb{R}^{2\times 2}$ has eigendecomposition $\Sigma_{\pi}=Q_{\pi}D_{\pi}Q_{\pi}^{T}$ with $D_{\pi}=\text{diag}\{\lambda_{1},\lambda_{2}\}$. We assume that $\Sigma_{\pi}$ is not a multiple of the identity. Construct the preconditioner $L=Q_{\pi}GD_{\pi}^{-1/2}G^{T}Q_{\pi}^{T}$ with the correct eigenvalues ($\epsilon=0$ in Assumption 3.1) but whose eigenvectors have been perturbed by an orthogonal matrix $G$ from those of $\Sigma_{\pi}$ by the angle $\arccos(1-\delta)$. It can be shown that the coefficient of $\delta^{4}$ in $\kappa_{L}$ is $(1/4)\times(l-2)^{2}$ where $l:=\lambda_{1}\lambda_{2}^{-1}+\lambda_{1}^{-1}\lambda_{2}$. So the more anisotropic $\Sigma_{\pi}$ is the more we are punished for having misaligned eigenvectors, as stated in the remark above. Figure 2 illustrates this fact. Each plot contains two contours: a blue one representing $\mathcal{N}(0,\Sigma_{\pi})$ and an orange one representing $\mathcal{N}(0,(LL^{T})^{-1})$. In both cases $G$ perturbs the eigenvectors by $\pi/4$, the angle between the semi-major axes of the contours is shown in the red arrows. In the first case we have $(\lambda_{1},\lambda_{2})=(2,1)$, in the second we have $(\lambda_{1},\lambda_{2})=(50,1)$ engendering a far smaller ‘overlap’ than in the first. The fact that $\kappa_{L}=O(\delta^{4})$ also shows that the $\delta$ dependency in Theorem 3 is tight.
Figure 2: Two pairs of contour plots, each representing $\mathcal{N}(0,\Sigma_{\pi})$ (blue) and $\mathcal{N}(0,(LL^{T})^{-1})$ (orange). The angle between the semi-major axes (red) is the same in either case, but the preconditioner in the right hand plot is worse due to the anisotropy of $\mathcal{N}(0,\Sigma_{\pi})$.
Assumption 3.1 and Assumption 3.2 require knowledge of each individual eigenvalue and eigenvector of $\nabla^{2}U$ across the entire state space, which may not always be available. We next provide more easily verifiable assumptions under which a similar result holds. This is achieved using results from matrix perturbation theory. The eigenvalues of $\nabla^{2}U$ can be controlled using only knowledge of the spectral norm by Weyl’s inequality (see Assumption 3.3). Similarly, using the Davis–Kahan theorem (e.g. Yu et al. (2015)) eigenvectors can be controlled by the spectral norm provided that an ‘eigengap’ condition holds (see Assumption 3.4). We therefore provide a second result under Assumptions 3.3 and 3.4 below.
Assumption 3.3.
There exists an $\epsilon>0$ such that $\|\nabla^{2}U(x)-LL^{T}\|\leq\sigma_{d}^{2}\epsilon$ for all $x\in\mathbb{R}^{d}$.
Assumption 3.4.
That $\gamma>0$ where
$\gamma:=\inf_{\begin{subarray}{c}i,j\in[d]\\ |i-j|=1\end{subarray}}|\sigma_{i}^{2}-\sigma_{j}^{2}|$
is the eigengap of $LL^{T}$.
Theorem 4.
Let $\Pi$ have potential $U$ satisfying Assumption 2.1. For a given preconditioner $L\in GL_{d}(\mathbb{R})$ for which Assumptions 3.3 and 3.4 hold, the condition number after preconditioning satisfies
$\kappa_{L}\leq(1+\epsilon)^{2}\left(1+\delta\sqrt{\sum_{i=1}^{d}\sigma_{i}^{2}\sum_{i=1}^{d}\sigma_{i}^{-2}}\right)^{4}$
where $\delta:=1-(1-2\gamma^{-1}\sigma_{d}^{-2}\epsilon)^{2}$.
For a proof see Appendix A.7. Based on this result, it might be tempting to arbitrarily increase the eigengap $\gamma$ or the least eigenvalue $\sigma_{d}$ of $L$. Note, however, that this may also cause $\epsilon$ to increase.
Remark.
Appendix A.6 shows that Assumption 3.3 implies Assumption 3.1. Similarly Assumption 3.3 and Assumption 3.4 combined imply Assumption 3.2 (Yu et al., 2015, Corollary 1). If the norm $\|\nabla^{2}U(x)-LL^{T}\|$ is difficult to compute the Frobenius norm can be used to form the upper bound since $\|\nabla^{2}U(x)-LL^{T}\|\leq\|\nabla^{2}U(x)-LL^{T}\|_{F}$.
Remark.
In the case of the additive Hessian an application of the triangle inequality gives
$\|A+B(x)-LL^{T}\|\leq\min\{\|A-LL^{T}\|+\|B(x)\|,\|B(x)-LL^{T}\|+\|A\|\}$
for all $x\in\mathbb{R}^{d}$. Therefore Assumption 3.3 suggests choosing $L=A^{1/2}$ if $A$ is ‘larger’ than $B(x)$ across $x\in\mathbb{R}^{d}$ and $L=B(x^{*})^{1/2}$ for some $x^{*}\in\mathbb{R}^{d}$ if $B(x)$ is larger, but does not exhibit large variations across $\mathbb{R}^{d}$.
The eigengap condition Assumption 3.4 is always satisfiable as we are free to choose $L$. It may not, however, be desirable. We therefore present a final bound on $\kappa_{L}$ that requires only Assumption 3.3.
Theorem 5.
Let $\Pi$ have potential $U$ satisfying Assumption 2.1. For a given preconditioner $L\in GL_{d}(\mathbb{R})$ for which Assumption 3.3 holds, the condition number after preconditioning satisfies
$\kappa_{L}\leq(1+\epsilon)\left(1+\frac{\sigma_{1}^{2}}{m}\epsilon\right)$
For a proof see Appendix A.8.
3.3 Linear preconditioning for multiplicative Hessians
A multiplicative Hessian has the form
$\nabla^{2}U(x)=X^{T}\Lambda(x)X$ (9)
where $X\in\mathbb{R}^{n\times d}$ for $n\geq d$ is a matrix whose rows are usually the rows of some dataset and $\Lambda(x)\in\mathbb{R}^{n\times n}$. An example of a model with a multiplicative Hessian is as follows
$\pi(\theta)\propto\exp\left(-\sum_{k=1}^{n}l_{y_{k}}(x_{k}^{T}\theta)-\frac{\lambda}{2}(\theta-\mu)^{T}X^{T}\Lambda X(\theta-\mu)\right)$ (10)
where $\{(y_{k},x_{k})\}_{k=1}^{n}$ are observations with $y_{k}\in\mathbb{R}$ and $x_{k}\in\mathbb{R}^{d}$ for $k\in[n]$, $X$ is a matrix with element in row $i$ and column $j$ equal to the $j$th element of $x_{i}$, and $\Lambda\in\mathbb{R}^{n\times n}$ is positive definite. Here $\ell_{y_{k}}$ denotes some loss associated with observation $k$. In the case $\ell_{y_{k}}$ is a negative log-likelihood, equation (10) therefore describes the posterior associated with a typical generalised linear model using the generalised g-prior of Hanson et al. (2014).
Without further assumptions we have the following:
Proposition 6.
A measure $\Pi$ whose potential $U$ has a multiplicative Hessian satisfies
$\frac{\sup_{x\in\mathbb{R}^{d}}\lambda_{n-d+1}(\Lambda(x))}{\inf_{x\in\mathbb{R}^{d}}\lambda_{d}(\Lambda(x)\kappa(X^{T}X)}\leq\kappa\leq\kappa(X^{T}X)\frac{\sup_{x\in\mathbb{R}^{d}}\lambda_{1}(\Lambda(x))}{\inf_{x\in\mathbb{R}^{d}}\lambda_{d}(\Lambda(x))}$
The proof in Appendix A.9 relies on an extension of Ostrowski’s theorem to rectangular matrices (Higham and Cheng, 1998a, Theorem 3.2). In many cases $\Lambda(x)$ will be diagonal and each eigenvalue $\lambda_{i}(\Lambda(x))$ will range between the same possible values $c$ and $C$ (this is the case for binary logistic regression using the g-prior, for example). In this instance a more precise statement about $\kappa$ can straightforwardly be made.
Assumption 3.5.
The Hessian of $\Pi$ is of multiplicative form with $\Lambda(x)$ diagonal, and there exists $c,C>0$ such that $\sup_{x\in\mathbb{R}^{d}}\lambda_{i}(\Lambda(x))=C$ and $\inf_{x\in\mathbb{R}^{d}}\lambda_{i}(\Lambda(x))=c$ for all $i\in[n]$.
Proposition 7.
A measure $\Pi$ for which Assumption 3.5 holds has a condition number
$\kappa=\frac{C}{c}\kappa(X^{T}X)$
A natural choice of preconditioner here is $L=(X^{T}X)^{1/2}$. Indeed $L=(X^{T}X)^{1/2}$ is proportional to the preconditioner suggested by Dalalyan (2017, section 6.2). For this we have the following.
Proposition 8.
Consider a measure $\Pi$ whose potential has a multiplicative Hessian. Then preconditioning with $L=(X^{T}X)^{1/2}$ gives
$\frac{\sup_{x\in\mathbb{R}^{d}}\lambda_{n-d+1}(\Lambda(x))}{\inf_{x\in\mathbb{R}^{d}}\lambda_{d}(\Lambda(x))}\leq\kappa_{L}\leq\frac{\sup_{x\in\mathbb{R}^{d}}\lambda_{1}(\Lambda(x))}{\inf_{x\in\mathbb{R}^{d}}\lambda_{d}(\Lambda(x))}$
For a proof see Appendix A.10. Under Assumption 3.5 the upper and lower bounds of Proposition 8 are equal, giving:
Corollary 9.
Consider a measure $\Pi$ for which Assumption 3.5 holds. Choosing $L=(X^{T}X)^{1/2}$ gives
$\kappa_{L}=\frac{C}{c}$
Hence the condition number is clearly reduced under preconditioning with $L=(X^{T}X)^{1/2}$. Another natural choice of preconditioner is $L=\nabla^{2}U(x^{*})^{1/2}=(X^{T}\Lambda(x^{*})X)^{1/2}$ for some $x^{*}\in\mathbb{R}^{d}$. For instance $x^{*}$ could be the mode of $\Pi$. Here we have the following
Proposition 10.
Consider a measure $\Pi$ whose potential has a multiplicative Hessian. Then choosing $L=(X^{T}\Lambda(x^{*})X)^{1/2}$ gives
$\kappa_{L}\leq\frac{\sup_{x\in\mathbb{R}^{d}}\lambda_{1}(\Lambda(x^{*})^{-1/2}\Lambda(x)\Lambda(x^{*})^{-1/2})}{\inf_{x\in\mathbb{R}^{d}}\lambda_{d}(\Lambda(x^{*})^{-1/2}\Lambda(x)\Lambda(x^{*})^{-1/2})}\leq\left(\frac{\sup_{x\in\mathbb{R}^{d}}\lambda_{1}(\Lambda(x))}{\inf_{x\in\mathbb{R}^{d}}\lambda_{d}(\Lambda(x))}\right)^{2}$
Note that the first upper bound is simply the condition number of $\tilde{\Pi}$, a measure whose potential has Hessian $\Lambda(x)$, after preconditioning with $L=\Lambda(x^{*})^{1/2}$. Therefore if $\tilde{\Pi}$ is preconditionable with $L=\Lambda(x^{*})^{1/2}$ then $\Pi$ is preconditionable with $L=(X^{T}\Lambda(x^{*})X)^{1/2}$. A proof of the proposition can be found in Appendix A.11.
We test the effectiveness of various preconditioners, including $L=(X^{T}X)^{1/2}$ and $L=(X^{T}\Lambda(x^{*})X)^{1/2}$ where $x^{*}\in\mathbb{R}^{d}$ is the mode, in the experiment in Section 4.3.
3.4 Tight condition number dependence of the spectral gap of random walk Metropolis
The condition number is of interest as it can be related to quantities such as the mixing and relaxation time of algorithms. Recent works (Dwivedi et al., 2018; Chen et al., 2020; Chewi et al., 2021; Lee et al., 2021; Andrieu et al., 2022) have established bounds on these quantities, which, in every case, rely polynomially on both the condition number and dimension (see Table 1). In particular Andrieu et al. (2022) give upper and lower bounds on the spectral gap of RWM which are tight in dimension, but only the lower bound is explicit in its dependence on the condition number. We posit the following assumption to make both bounds explicit.
Assumption 3.6.
That $\Pi$ satisfies Assumption 2.1 and there exists an $\epsilon>0$ such that $\|\nabla^{2}U(x)-\nabla^{2}U(y)\|\leq m\epsilon$ for all $x,y\in\mathbb{R}^{d}$.
Theorem 11.
Let $\Pi$ have potential $U$ satisfying Assumption 2.1 with condition number $\kappa=M/m\geq 1$. If $\Pi$ also satisfies Assumption 3.6 then the spectral gap $\gamma_{\kappa}$ of the random walk Metropolis with proposal variance $\sigma^{2}I$ such that $\sigma^{2}:=\xi/(Md)$ for any $\xi>0$ satisfies
$C\xi\exp(-2\xi)\cdot\frac{1}{\kappa}\cdot\frac{1}{d}\leq\gamma_{\kappa}\leq(1+2\epsilon)\frac{\xi}{2}\cdot\frac{1}{\kappa}\cdot\frac{1}{d}$
where $C=1.972\times 10^{-4}$
For a proof see Appendix A.12. Both bounds presented above are $O(\kappa^{-1}d^{-1})$, which implies that the relaxation time $1/\gamma_{\kappa}$ is precisely linear in $\kappa$. The lower bound is as originally presented by Andrieu et al. (2022, Theorem 1), and the choice $\sigma^{2}=\xi/(Md)$ is as recommended in that work to ensure tight $O(d^{-1})$ dependence. As the authors remark, the constant $C$ can possibly be made a few orders of magnitude larger.
Theorem 11 can be combined with the results of Section 3.2 to guarantee that the spectral gap increases under linear preconditioning.
Corollary 12.
Let $\Pi$ have potential $U$ satisfying Assumption 2.1 with condition number $\kappa=M/m\geq 1$. Assume that $\Pi$ also satisfies Assumption 3.6 with constant $\epsilon^{\prime}>0$. Using a preconditioner $L\in GL_{d}(\mathbb{R})$ satisfying Assumption 3.3 with constant $\epsilon>0$ ensures that the spectral gap $\gamma_{\kappa}$ of the RWM with proposal variance $\sigma^{2}I$ such that $\sigma^{2}:=\xi/(Md)$ for any $\xi>0$ increases under preconditioning whenever
$\kappa\geq\frac{1}{2}C^{-1}\exp(2\xi)(1+2\epsilon^{\prime})(1+\epsilon)\left(1+\frac{\sigma_{1}(L)^{2}}{m}\epsilon\right)$
where $C=1.972\times 10^{-4}$, $\sigma_{d}(L)$ is the least singular value of $L$.
Corollary 12 uses Theorem 11 and Theorem 5. A proof can be found in A.13. We could have stated a similar result using Theorem 3 or Theorem 4 in the place of Theorem 5 which would change the necessary lower bound on $\kappa$ in the above result.
3.5 Common choices of linear preconditioner
In this section we offer an overview of existing choices of linear preconditioners. Before describing the contents of the matrices, we note that practitioners must first choose their sparsity structure. The most popular choices are either fully dense and diagonal. A diagonal preconditioner offers clear computational advantages, and is therefore preferable in many settings. In other settings model-specific knowledge can be exploited to design preconditioners with a particular sparsity structure (for example, see Wallin and Bolin (2018)).
Using $L=\Sigma_{\pi}^{-1/2}$ is a popular choice. Its use extends at least as far back as Haario et al. (2001), with off-the-shelf options presented in many major MCMC software packages. See Section 3.5.2 for a theoretical examination of its effectiveness and Sections 4.2 and 4.3 for a numerical study.
3.5.1 The Fisher matrix
Titsias (2023) suggests using $L\propto\mathcal{I}^{1/2}$ where $\mathcal{I}:=\mathbb{E}_{\Pi}[\nabla U(x)\nabla U(x)^{T}]$ is called the Fisher matrix. This choice of $L$ maximises the expected squared jump distance of the unadjusted Langevin sampler, which is simply the Euler-Maruyama discretisation of (1). A straightforward integration by parts shows that $\mathcal{I}$ can also be written $\mathbb{E}_{\Pi}[\nabla^{2}U(x)]$, which highlights its relationship with $\nabla^{2}U$. Proposition 13 shows that if a choice of $L$ satisfying Assumption 3.3 exists, then the alternative choice of preconditioner $\mathcal{I}^{1/2}$ will also be valid.
Proposition 13.
Let $\Pi$ have potential $U\in C^{2}$ and assume there exists a preconditioner $L\in GL_{d}(\mathbb{R})$ satisfying Assumption 3.3 for some $\epsilon>0$. Then $\|\nabla^{2}U(x)-\mathcal{I}\|\leq 2\sigma_{d}^{2}\epsilon$ for all $x\in\mathbb{R}^{d}$ where $\sigma_{d}^{2}$ is the least eigenvalue of $LL^{T}$.
Proof.
The assumption has that $\|\nabla^{2}U(x)-LL^{T}\|\leq\sigma_{d}^{2}\epsilon$ for all $x\in\mathbb{R}^{d}$. Therefore
$\|\nabla^{2}U(x)-\mathcal{I}\|$ $=\|\nabla^{2}U(x)-LL^{T}+LL^{T}-\mathcal{I}\|$
$\leq\sigma_{d}^{2}\epsilon+\left\|\mathbb{E}_{\Pi}[\nabla^{2}U(X)-LL^{T}]\right\|$
$\leq\sigma_{d}^{2}\epsilon+\mathbb{E}_{\Pi}[\|\nabla^{2}U(X)-LL^{T}\|]$
$\leq 2\sigma_{d}^{2}\epsilon$
where the penultimate line is due to Jensen’s inequality. ∎∎
Corollary 14.
Consider $\Pi$ satisfying Assumption 2.1 with a potential $U\in C^{2}$ and a preconditioner $L\in GL_{d}(\mathbb{R})$ satisfying Assumption 3.3 for some $\epsilon>0$. Then
$\kappa_{\mathcal{I}^{\frac{1}{2}}}\leq(1+2\epsilon)\left(1+\frac{\sigma_{1}^{2}}{m}2\epsilon\right)$
where $\sigma_{1}^{2}$ is the greatest eigenvalue of $LL^{T}$.
The proof is a direct application of Theorem 5. Section 4.3 offers a numerical test of $L=\Sigma_{\pi}^{-1/2}$ and $L=\mathcal{I}^{1/2}$.
Another commonly used choice of linear preconditioner is to set $L=\nabla^{2}U(x^{*})^{1/2}$ where $x^{*}$ is chosen to make $\nabla^{2}U(x^{*})$ sufficiently representative of $\nabla^{2}U(x)$ everywhere (e.g. the mode). This choice avoids computing expectations with respect to $\Pi$. We test this choice of preconditioner in Section 4.3.
3.5.2 The target covariance
Preconditioning with an estimate of the target covariance is a popular strategy. This is not without theoretical justification: here we look at how the spectral gap of the Ornstein–Uhlenbeck (O-U) process changes under preconditioning. The preconditioned O-U process is an instance of the preconditioned Langevin diffusion which is driven by the stochastic differential equation:
$dX_{t}=-\frac{1}{2}(LL^{T})^{-1}\nabla U(X_{t})dt+L^{-1}dB_{t},$ (11)
where $B_{t}$ is a Brownian motion. The preconditioned O-U process is found by setting $\Pi=\mathcal{N}(0,\Sigma_{\pi})$, giving
$dX_{t}=-\frac{1}{2}(LL^{T})^{-1}\Sigma_{\pi}^{-1}X_{t}dt+L^{-1}dB_{t}.$
Since our practical goal is to simulate this process on a computer, we must take care when interpreting results about its continuous time formulation. Therefore we subject ourselves to the condition $\left|\det(-L^{-1}L^{-T}\Sigma_{\pi}^{-1})\right|=1$. If this were not done we could choose $L^{\prime}:=s^{-1}L$ for increasingly large $s>0$, which would arbitrarily increase the rate of convergence of the continuous time process, but destabilise the discretised process for any fixed numerical integrator step-size. The below result is most likely well-known, but we were unable to find a reference, and so provide it here for completeness.
Proposition 15.
The preconditioner $L=\Sigma_{\pi}^{-1/2}$ maximises the spectral gap of the preconditioned O-U process subject to $\left|\det(-L^{-1}L^{-T}\Sigma_{\pi}^{-1})\right|=1$.
Proof of Proposition 15 can be found in Appendix A.14. Alrachid et al. (2018) also identify the spectral gap of the preconditioned O-U process, although since they focus on the continuous time dynamics, they do not identify an optimal preconditioner (with respect to the spectral gap).
In section 3.2 we characterised the performance of a preconditioner $L$ by how much the Hessian $\nabla^{2}U(x)$ varies around $LL^{T}$. Here we do the same for $L=\Sigma_{\pi}^{-1/2}$.
Proposition 16.
Let $\Pi$ be a measure with potential $U$, covariance $\Sigma_{\pi}\in\mathbb{R}^{d\times d}$, mode $x^{*}\in\mathbb{R}^{d}$, and expectation $\mu_{\pi}\in\mathbb{R}^{d}$. Assume that there exist positive definite matrices $\Delta_{+},\Delta_{-}\in\mathbb{R}^{d\times d}$ such that $\Delta_{-}\preceq\nabla^{2}U(x)\preceq\Delta_{+}$ for all $x\in\mathbb{R}^{d}$ and that $01-(x^{*}-\mu_{\pi})^{T}\Delta_{+}(x^{*}-\mu_{\pi})>$ 0. Then $P_{-}\preceq\Sigma_{\pi}^{-1}\preceq P_{+}$ where
$P_{+}$ $=c^{-1}\left(\mathbf{I}_{d}+\left(1-\text{\emph{Tr}}(D_{+})\right)^{-1}D_{+}\right)\Delta_{+}$
$P_{-}$ $=c\left(\mathbf{I}_{d}+\left(1-\text{\emph{Tr}}(D_{-})\right)^{-1}D_{-}\right)\Delta_{-}$
with $D_{\pm}=\Delta_{\pm}(x^{*}-\mu_{\pi})(x^{*}-\mu_{\pi})^{T}$ and $c:=\sqrt{\det\Delta_{-}{\det\Delta_{+}^{-1}}}\leq 1$, and in addition
$\|\nabla^{2}U(x)-\Sigma_{\pi}^{-1}\|\leq\max\{\|\Delta_{+}-P_{-}\|,\|P_{+}-\Delta_{-}\|\}$
for all $x\in\mathbb{R}^{d}$.
For a proof see Appendix A.15. Proposition 16 allows us to localise the covariance in terms of the parameters of the target distribution.
The intuition from Section 3.2 suggests that Hessians which exhibit small variations across the state space are preconditionable. In this scenario we would have $\Delta_{+}\approx\Delta_{-}$ and hence that $c\approx 1$. Proposition 16 then suggests that so long as the distance between the mean and the mode is not too great, $\|\nabla^{2}U(x)-\Sigma_{\pi}^{-1}\|$ is small. In summary, if $\Pi$ is preconditionable, and if the mean is close to the mode, then preconditioning with $L=\Sigma_{\pi}^{-1/2}$ is sensible.
Recall that a potential with additive Hessian satisfies $\nabla^{2}U(x)=A+B(x)$ where $A,B(x)\in\mathbb{R}^{d\times d}$ are symmetric. See Section 3.2 for models whose potentials have additive Hessians. In this case $\Delta_{-}$ and $\Delta_{+}$ will be generated by variations in $B(x)$. Therefore, given Proposition 16, a tighter localisation of $B(x)$ gives a tighter localisation of $\nabla^{2}U(x)$ around the inverse covariance:
Corollary 17.
Let $\Pi$ be a measure with potential $U$, covariance $\Sigma_{\pi}\in\mathbb{R}^{d\times d}$, mode $x^{*}\in\mathbb{R}^{d}$, and expectation $\mu_{\pi}\in\mathbb{R}^{d}$. If the Hessian of $U$ is of the form $\nabla^{2}U(x)=A+B(x)$ with $\|B(x)\|\leq\epsilon$ and $\epsilon\mathbf{I}_{d}\prec A$ for some $\epsilon>0$, then $01-(x^{*}-\mu_{\pi})^{T}(A+\epsilon\mathbf{I}_{d})(x^{*}-\mu_{\pi})>$ 0 ensures that
$\|\nabla^{2}U(x)-\Sigma_{\pi}^{-1}\|\leq(c^{-1}+1)\epsilon+(c^{-1}-1)\|A\|+\max\{\|\tilde{P}_{-}\|,\|\tilde{P}_{+}\|\},$
where
$\tilde{P}_{+}$ $=c^{-1}\left(1-\text{\emph{Tr}}(\tilde{D}_{+})\right)^{-1}\tilde{D}_{+}(A+\epsilon\mathbf{I}_{d}),$
$\tilde{P}_{-}$ $=c\left(1-\text{\emph{Tr}}(\tilde{D}_{-})\right)^{-1}\tilde{D}_{-}(A-\epsilon\mathbf{I}_{d}),$
with $\tilde{D}_{\pm}=(A\pm\epsilon\mathbf{I}_{d})(x^{*}-\mu_{\pi})(x^{*}-\mu_{\pi})^{T}$ and c:=$\sqrt{{\det(A-\epsilon\mathbf{I}_{d})}{\det(A+\epsilon\mathbf{I}_{d})^{-1}}}\leq 1$.
Proof can be found in Appendix A.16. In Section 4.2 we look at the difference in performance between the preconditioners $L=A^{1/2}$, $L=\Sigma_{\pi}^{-1/2}$ and $L=\mathbf{I}_{d}$.
3.5.3 Counterproductive diagonal preconditioning
The appeal of diagonal preconditioning is motivated by the promise of reducing the condition number in $O(d)$ computational cost. The practice has a strong tradition in numerical linear algebra, where diagonal preconditioning is commonplace. Therefore choosing $L=\text{diag}(\hat{\Sigma}_{\pi})^{-1/2}$ for some estimate $\hat{\Sigma}_{\pi}$ of the target covariance has become a common practice since it is viewed as computationally cheap, and it is assumed that it will offer an improvement on no preconditioning at all. For instance the developers of Stan also offer the option of diagonal preconditioning with the target covariance. Such an option is also offered in the TensorFlow Probability library (Abadi et al., 2015). However, contrary to its appeal, below we provide examples of diagonal preconditioning that actually increase the condition number.
Even when the target is Gaussian, diagonal preconditioning can lead to a worse-performing algorithm. Noting that $\kappa_{L}=\|\text{diag}(\Sigma_{\pi})^{1/2}\Sigma_{\pi}^{-1}\text{diag}(\Sigma_{\pi})^{1/2}\|\|\text{diag}(\Sigma_{\pi})^{-1/2}\Sigma_{\pi}\text{diag}(\Sigma_{\pi})^{-1/2}\|=\|C_{\pi}^{-1}\|\|C_{\pi}\|$ where $C_{\pi}\in\mathbb{R}^{d\times d}$ is the correlation matrix associated with $\Sigma_{\pi}$, it suffices to find covariance matrices $\Sigma_{\pi}$ whose spectral condition number increases when we look at their correlation matrices. The matrix below is such an example.
$\Sigma_{\pi}=\begin{pmatrix}21.5&5.7&18.7&4.5&6.9\\ \star&2.0&4.9&1.2&2.1\\ \star&\star&16.3&3.9&5.7\\ \star&\star&\star&1.4&1.4\\ \star&\star&\star&\star&2.9\end{pmatrix}\implies\kappa\approx 4.4\times 10^{3},\kappa_{L}\approx 8.1\times 10^{3}$ (12)
The condition number increases by a substantial amount, even though we have perfect knowledge of the target covariance. See Section 4.1 for an empirical analysis of RWM on a $\mathcal{N}(0,\Sigma_{\pi})$ target (with $\Sigma_{\pi}$ as above) after diagonal and dense preconditioning.
In a practical scenario we would have to expend computational effort to construct $\text{diag}(\hat{\Sigma}_{\pi})^{-1/2}$ and so for a target such as the one described here, the effort would be for nothing. In general targets with covariance matrices whose associated correlations are far from being diagonally dominant will be least amenable to diagonal preconditioning.
In fact, it is a simple corollary to Theorem 2 of Qu et al. (2020) that
$\kappa(C_{\pi})\leq d\left(\frac{1-\alpha}{1+\alpha}\right)^{2}\frac{\min_{i\in[d]}(\Sigma_{\pi})_{ii}}{\max_{i\in[d]}(\Sigma_{\pi})_{ii}}\kappa(\Sigma_{\pi})$
where $\kappa(.)$ is the spectral condition number, and $C_{\pi}$ is $\alpha$-diagonally dominant (i.e. there exists an $\alpha>0$ such that $1\geq\alpha\sum_{j\neq i}^{d}|(C_{\pi})_{ij}|$ for all $i\in[d]$). So $\kappa(C_{\pi})$ will exist in a larger range for $C_{\pi}$ which are far from diagonal dominance.
4 Experiments
4.1 Counterproductive Diagonal Preconditioning
This experiment illustrates the phenomenon described in section 3.5.3: namely that there exist Gaussian targets in the form $\mathcal{N}(0,\Sigma_{\pi})$ such that preconditioning with $L=\text{diag}(\Sigma_{\pi})^{-1/2}$ increases the condition number.
We compare the performance of three RWM algorithms: one with no preconditioning, one with dense preconditioning ($L=\Sigma_{\pi}^{-1/2}$), and one with diagonal preconditioning ($L=\text{diag}(\Sigma_{\pi})^{-1/2}$). Each chain targets $\mathcal{N}(0,\Sigma_{\pi})$ with $\Sigma_{\pi}$ as in (12) and is initialised at equilibrium. Proposals take the standard RWM form $X^{\prime}=X+\sigma L^{-1}\xi$ with $\xi\sim\mathcal{N}(0,\mathbf{I}_{5})$ and $\sigma=2.38/\sqrt{d}$ as recommended by Roberts and Rosenthal (2001). We run each chain from each algorithm 100 times at 10,000 iterations per chain. For each chain we compute the effective sample size (ESS) in each dimension using the effectiveSize function from the coda package (Plummer et al., 2006). The ESS is an estimator which measures the amount of independent samples one would need to achieve an empirical average with equivalent variance to the one computed from the chain.
Figure 3: $\log\text{ESS}$ of 100 runs at 10,000 iterations per run of RWM with dense, diagonal, and zero preconditioning. Each algorithm is started in equilibrium and targets $\mathcal{N}(0,\Sigma_{\pi})$ with $\Sigma_{\pi}$ as in (12).
As can be seen in Figure 3 the ESS’s of the RWM chain with no preconditioning are clearly larger in dimensions 2, 4, and 5 from the diagonally preconditioned chain. This is despite the diagonal preconditioner being formed with perfect knowledge of the target covariance. As is expected the dense preconditioner performs the best since the target effectively becomes a standard Gaussian.
4.2 Preconditioning the additive Hessian using the covariance
One probabilistic model with additive Hessian structure is a Bayesian regression with hyperbolic prior. It is well known that the Laplace prior $\beta\mapsto 2^{-1}\lambda\exp(-\lambda\|\beta\|_{1})$ for $\beta\in\mathbb{R}^{d}$ and $\lambda>0$ imposes the same sparsity in the maximum a posteriori estimates as would regularisation with the LASSO (Tibshirani, 1996) since it concentrates sharply around $\beta=0$. More generally, Laplacian tails can be motivated by results concerning the contraction of predictors around their true values. As demonstrated in the discussion following Theorem 8 in Castillo et al. (2015), heavy-tailed priors such as the Laplacian achieve good rates of contraction. Simply using a Laplacian prior would violate the $M$-smoothness assumption due to the sharpness at $\beta=0$. The hyperbolic prior, however, is a smooth distribution with Laplacian tails. We assume that $Y=X\beta+\epsilon$ where $\epsilon\sim\mathcal{N}(0,\sigma^{2}\mathbf{I}_{n})$ for $\sigma^{2}>0$ known and $n>d$. We also assume that the columns of $X\in\mathbb{R}^{n\times d}$ are standardised to have variance 1. Because of this it is reasonable to use the same scale in each dimension of the prior. The resulting posterior has a potential of the form
$U(\beta)=\frac{1}{2\sigma^{2}}\|Y-X\beta\|^{2}+\lambda\sum_{i=1}^{d}\sqrt{1+\beta_{i}^{2}},$ (13)
which implies
$\nabla^{2}U(\beta)=\frac{1}{\sigma^{2}}X^{T}X+\lambda D(\beta),$ (14)
where $D(\beta)=\text{diag}\{(1+\beta_{i}^{2})^{-3/2}:i\in[d]\}$. The Hessian is therefore additive with $A=\sigma^{-2}X^{T}X$ and $B(\beta)=\lambda D(\beta)$.
As can be seen from (13)-(14), $U$ satisfies Assumption 2.1 with $m=\sigma^{-2}\sigma_{d}(X^{T}X)$ and $M=\sigma^{-2}\|X^{T}X\|+\lambda$. Preconditioning with $L=(\sigma^{-2}X^{T}X)^{1/2}$ gives $\kappa_{L}=1+\sigma_{d}(X^{T}X)^{-1}\lambda\sigma^{2}\leq\kappa$. In this case the distance between $LL^{T}$ and the Hessian can be bounded using $\|\nabla^{2}U(\beta)-LL^{T}\|\leq\lambda$. We can therefore apply Proposition 17 by setting $\epsilon=\lambda$, implying that when $\lambda$ is small then preconditioning with either $L=(\sigma^{-2}X^{T}X)^{1/2}$ or $L=\Sigma_{\pi}^{-1/2}$ should improve the efficiency of the sampler, so long as the distance between the mean and the mode is not too large.
In the following experiment we run MALA chains on target distributions with $L=(\sigma^{-2}X^{T}X)^{1/2}$, $L=\Sigma_{\pi}^{-1/2}$, and $L=\mathbf{I}_{d}$. We set $d\in\{2,5,10,20,100\}$ and $n=\{1,5,20\}\times d$ for each value of $d$. At each combination of $n$ and $d$ we run 15 chains for each preconditioner. Each chain is composed by initialising at $\beta=(X^{T}X)^{-1}X^{T}Y$ and taking $10^{4}$ samples to equilibrate. We initialise the step size at $d^{-1/6}$ and adapt it along the course of the chain seeking an optimal acceptance rate of $0.574$ according to the results of Roberts and Rosenthal (2001). We then continue the chain with preconditioning and a fixed step size of $d^{-1/6}$ for a further $10^{4}$ samples, over which we measure the ESS of each dimension. To construct $L=\Sigma_{\pi}^{-1/2}$ we simply use the empirical covariance of the first $10^{4}$ samples.
For the model parameters we set $\lambda=\sqrt{n}/d$ using the lower bound of Castillo et al. (2015). Every element of $X$ is an independent standard normal random variable, and $Y$ is generated by sampling $\beta_{0}$ from the prior and setting $Y=X\beta_{0}+\epsilon$ with $\epsilon\sim\mathcal{N}(0,\mathbf{I}_{d})$. Therefore $\sigma=1$.
The boxplots in Figure 4 show the median ESSs for each run. The figure demonstrates that in the $n/d\in\{5,20\}$ cases preconditioning with $L=\Sigma_{\pi}^{-1/2}$ is just as good as preconditioning with $L=A^{1/2}$ where $A=\sigma^{-1}X^{T}X$. For $n/d=1$ the results are mixed: for instance in the $(d,n)=(100,100)$ configuration the first $10^{4}$ iterations of the MALA chain mixed poorly, offering a poor estimate of $\Sigma_{\pi}$. The performance suffered heavily if no preconditioner was applied.
Figure 4: Boxplots of the medians of the ESSs across configurations of $(n,d)$ with different preconditioners on the Bayesian linear regression with a Hyperbolic prior. The leftmost boxplot in each grouping corresponds to preconditioning with $L=\sigma(X^{T}X)^{1/2}$, the middle boxplot has $L=\Sigma_{\pi}^{-1/2}$, the rightmost has $L=\mathbf{I}_{d}$.
4.3 Preconditioning the multiplicative Hessian
To study preconditioning under the multiplicative Hessian structure we consider a Bayesian binomial regression with a generalised $g$-prior (Sabanés Bové and Held, 2010, Section 2.1) (Held and Sauter, 2017, Section 2.2). The generalised $g$-prior is an extension of the classical $g$-prior to generalised linear models that have dispersion parameters of the form $\phi_{i}:=\phi w_{i}^{-1}$ for $i\in[n]$ and known weights $w_{i}\in\mathbb{R}^{+}$. It is motivated by constructing an ‘imaginary sample’ of responses $y_{0}=h(0)\mathds{1}_{n}\in\mathbb{R}^{n}$ from a generalised linear model with inverse link function $h(.)$ and design matrix $X\in\mathbb{R}^{n\times d}$. Assigning the parameter vector $\beta\in\mathbb{R}^{d}$ a flat prior, it is observed that as $n\to\infty$ the posterior distribution of $\beta$ in this construction tends to $\mathcal{N}_{d}(0,g\phi c(X^{T}WX)^{-1})$ where $W=\text{diag}\{w_{i}:i\in[n]\}$, $g$ and $\phi$ are hyperparameters, and $c$ is a model-specific constant. See Sabanés Bové and Held (2010, Section 2.1) for a more thorough exposition.
We follow the advice given by Sabanés Bové and Held (2010, section 2.1) and Held and Sauter (2017, section 2.2) by setting $w_{i}=m_{i}$ for all $i\in[n]$. Using a logistic link gives a posterior with potential
$U(\beta)=\sum_{i=1}^{n}\left(w_{i}\left((1-Y_{i})X_{i}^{T}\beta+\log(1+\exp(-X_{i}^{T}\beta))\right)\right)+(g\phi c)^{-1}\beta^{T}X^{T}WX\beta$
with Hessian $\nabla^{2}U(\beta)=X^{T}\Lambda(\beta)X$, where
$\Lambda(\beta):=W\text{diag}\{\exp(X_{i}^{T}\beta)(1+\exp(X_{i}^{T}\beta))^{-2}+(g\phi c)^{-1}:i\in[n]\}$
The potential $U$ therefore satisfies Assumption 2.1 with $M=(0.25+(g\phi c)^{-1})w_{\text{max}}\|X^{T}X\|$ and $m=(g\phi c)^{-1}w_{\text{min}}\sigma_{d}(X^{T}X)$, where $w_{\text{max}}:=\max_{i}w_{i}$ and $w_{\text{min}}:=\min_{i}w_{i}$. We choose $g$ and $\phi$ such that $(g\phi c)^{-1}=\lambda n^{-1}$, where $\lambda=0.01$.
We examine the effectiveness of preconditioning with $L\in\{\Sigma_{\pi}^{-1/2},\mathcal{I}^{1/2},\nabla^{2}U(\beta^{*})^{1/2},\mathbf{I}_{d},(n^{-1}X^{T}X)^{1/2}\}$ where $\Sigma_{\pi}$ is the covariance of the posterior, $\mathcal{I}$ is the ‘Fisher matrix’ of Titsias (2023), $\beta^{*}$ is the mode, and $L=(n^{-1}X^{T}X)^{1/2}$ is the preconditioner used in Dalalyan (2017, Section 6.2). When $L\in\{\mathbf{I}_{d},(n^{-1}X^{T}X)^{1/2},\nabla^{2}U(\beta^{*})^{1/2}\}$ the condition numbers are given by
$L=\mathbf{I}_{d}$ $\Rightarrow\kappa_{L}=\kappa=\frac{\frac{n}{4}+\lambda}{\lambda}\frac{w_{\text{max}}}{w_{\text{min}}}\kappa(X^{T}X)$
$L=(n^{-1}X^{T}X)^{\frac{1}{2}}$ $\Rightarrow\kappa_{L}=\frac{\frac{n}{4}+\lambda}{\lambda}\frac{w_{\text{max}}}{w_{\text{min}}}$
$L=\nabla^{2}U(\beta^{*})^{\frac{1}{2}}$ $\Rightarrow\kappa_{L}=\frac{\frac{n}{4}+\lambda}{\lambda}\frac{n\max_{i}p_{i}^{*}(1-p_{i}^{*})+\lambda}{n\min_{i}p_{i}^{*}(1-p_{i}^{*})+\lambda}$
where $p_{i}^{*}:=(1+\exp(-X_{i}^{T}\beta^{*}))^{-1}$. This suggests that $L=\nabla^{2}U(\beta^{*})^{1/2}$ offers an increase in efficiency over $L=(n^{-1}X^{T}X)^{1/2}$ for $w_{\text{max}}/w_{\text{min}}$ large.
4.3.1 Experimental setup and results
We run RWM chains with the preconditioners described above for $d\in\{2,5,10,20\}$ and $n=5d$. We generate the design matrix $X\in\mathbb{R}^{n\times d}$ with $X=G+M$ where $G_{ij}\sim\mathcal{N}(0,1)$ independently and $M_{ij}=\mu$ for all $i\in[n],j\in[d]$. We set $\mu\in\{0,5,50,200\}$ to arbitrarily worsen the conditioning of the model, as it can be shown that
$\kappa(X^{T}X)\geq\frac{\sum_{k=1}^{n}(G_{k1}+\mu)^{2}}{\frac{1}{2}\sum_{k=1}^{n}(G_{k1}-G_{k2})^{2}}.$
We set $w_{i}=i^{2}$ for $i\in[n]$ and generate the responses using $Y_{i}=S_{i}/w_{i}$ with $S_{i}\sim\text{Bin}(w_{i},(1+\exp(-X_{i}^{T}\beta_{0}))^{-1})$ for $\beta_{0}\sim\mathcal{N}(0,\mathbf{I}_{d})$. We use gradient descent on $U$ which we precondition with $L=(n^{-1}X^{T}X)^{1/2}$ to find the mode $\beta^{*}$.
We approximate $\Sigma_{\pi}$ and $\mathcal{I}$ in two different ways. We either construct them using ergodic averages generated by unpreconditioned RWM for $10^{4}$ iterations, or we run an $L=\nabla^{2}U(\beta^{*})^{1/2}$ preconditioned RWM for $10^{5}$ iterations, from which we calculate the same ergodic averages. At each combination of $d$ and $\mu$ we run 15 chains for each preconditioner. Each chain is composed by initialising at $\beta\sim\mathcal{N}(0,(n^{-1}X^{T}X)^{-1})$ and taking $10^{4}$ samples to equilibrate. In each of these initial chains we initialise the step size at $2.38/d^{1/2}$ and adapt it along the course of the trajectory seeking an optimal acceptance rate of $0.234$ according to the results of Roberts and Rosenthal (2001). We then continue the chain with preconditioning and a fixed step size of $2.38/d^{1/2}$ for a further $10^{4}$ samples, over which we measure the ESS of each dimension. The median ESSs in the $\mu\in\{0,200\}$ cases are plotted in Figure 5.
Figure 5: Boxplots of the logarithms of the medians of the ESSs across combinations of $(d,\mu)$. The ESSs are taken from RWM runs with the specified preconditioning strategy on a binomial regression target with the generalised $g$-prior.
The preconditioning strategies are detailed in the legend. ‘covariance’ and ‘covarianceII’ refer to the runs preconditioned with $L=\Sigma_{\pi}^{-1/2}$ where the covariance is estimated over $10^{4}$ and $10^{5}$ samples respectively. The same is the case for ‘Fisher’ and ‘FisherII’. ‘sq_root_Sigma_X’ refers to the runs made with $L=(n^{-1}X^{T}X)^{1/2}$. ‘mode’ refers to preconditioning with $L=\nabla^{2}U(\beta^{*})^{1/2}$.
As predicted, preconditioning with the Hessian at the mode does offer a benefit over preconditioning with $L=(n^{-1}X^{T}X)^{1/2}$. Preconditioning with the covariance when it is estimated over a larger, better quality sample (‘covarianceII’) is one of the best performing strategies, whereas preconditioning with the covariance estimated over the smaller sample (‘covariance’) suffers with dimension and ill-conditioning of the model. This is clearly due to the reduction in quality of the covariance estimate. This disparity in performance is contrasted with the difference between the ‘Fisher’ and ‘FisherII’ cases, which is very slight.
5 Discussion
Our work offers many interesting extensions, which we now discuss in turn.
5.1 Alternative condition numbers and refinements
Alternative, problem specific condition numbers have been defined by various parties. For instance where $\Sigma_{\pi}$ is the covariance of $\Pi$ with spectrum $\sigma^{2}_{1}\geq\sigma^{2}_{2}\geq...\geq\sigma^{2}_{d}$ Langmore et al. (2020) suggest using
$\left(\sum_{i=1}^{d}\left(\frac{\sigma_{1}}{\sigma_{i}}\right)^{4}\right)^{\frac{1}{4}}$
as a condition number. Under specifications on the step size of the algorithm, such a quantity is shown to be proportional to the number of leapfrog steps needed to achieve a stable acceptance rate in HMC.
The condition number as defined in (4) encodes the difficulty of sampling from $\Pi$, but it does not capture additional information we might have about $\Pi$ which might ameliorate the sampling efficiency. For instance, if we knew that there existed positive definite $A_{-},A_{+}\in\mathbb{R}^{d\times d}$ such that
$A_{-}\preceq\nabla^{2}U(x)\preceq A_{+},$
then we could precondition with $L=A_{-}^{1/2}$ achieving $\kappa_{L}=\lambda_{1}(A_{-}^{-1}A_{+})$ over $\kappa=\lambda_{1}(A_{+})/\lambda_{d}(A_{-})$. Therefore defining the condition number as $\lambda_{1}(A_{-}^{-1}A_{+})$ encodes the difficulty of sampling given all the information at hand. See, for instance, Safaryan et al. (2021, Section 2.3), Saumard and Wellner (2014, Definition 2.9) or Hillion et al. (2019, Definition 1) for similarly motivated definitions.
5.2 Nonlinear preconditioning
A natural extension is to broaden the class of preconditioners to include nonlinear transformations. At present nonlinear preconditioning can be seen in the form of normalizing flows (Gabrié et al., 2022; Hoffman et al., 2019) and measure transport (Parno and Marzouk, 2018). Less computationally intensive transformations are considered by Johnson and Geyer (2012) and Yang et al. (2022) in order to sample from heavy-tailed distributions.
We note that to identify a transformation $g:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that the pushforward of $\Pi$ under $g$ has condition number 1 is to solve the equation
$U(x)+\log|\det J(g(x))|=\frac{1}{2}\|g(x)\|^{2}$
where $J(g(x))$ is the Jacobian of $g$ at $x\in\mathbb{R}^{d}$. This is an instance of the Monge-Ampère equation, which is well studied in optimal transport (Peyré and Cuturi, 2020, remark 2.25). Solvers of the Monge-Ampère exist in the literature, see Benamou et al. (2014, 2016b). Contextualising the existing analysis of the Monge-Ampère and its solvers within MCMC is a potentially fruitful line of inquiry.
There exist classes of algorithms that are equivalent to transforming existing sampling algorithms under nonlinear transformations. These include the Riemannian manifold algorithms of Girolami and Calderhead (2011) (see also Patterson and Teh (2013); Lan et al. (2015); Livingstone (2021)) and the algorithms derived from mirror descent (Nemirovski and Yudin, 1983) such as those seen in Hsieh et al. (2018); Zhang et al. (2020); Chewi et al. (2020). That the algorithms derived from mirror descent are equivalent to a nonlinearly preconditioned sampling scheme is evident in their construction. For the Riemannian manifold samplers, one can show, for instance, that the Langevin diffusion
$dY_{t}=\frac{1}{2}\nabla\log\tilde{\pi}(Y_{t})dt+dB_{t}$
under diffeomorphism $f(Y)=X$ transforms into the following SDE
$\begin{split}dX_{t}&=\frac{1}{2}G(X_{t})^{-1}\nabla\log\pi(X_{t})dt+\Gamma(X_{t})dt+G(X_{t})^{-\frac{1}{2}}dB_{t}\\ \Gamma_{i}(X_{t})&=\frac{1}{2}\sum_{j=1}^{d}\frac{\partial}{\partial x_{j}}\left(G(X_{t})_{ij}^{-1}\right)\end{split}$ (15)
with $G(x)^{-1}=J(g(x))^{-1}J(g(x))^{-T}$ where $g$ is the inverse of $f$ and $\pi(x)=\tilde{\pi}(y)|\det J(f(y))^{-1}|$, see Zhang et al. (2023a) for a formal statement and proof. Xifara et al. (2014); Livingstone and Girolami (2014) show that the SDE in 15 is the Langevin diffusion on the Riemannian manifold with metric $G(x)\in\mathbb{R}^{d\times d}$, and therefore the same diffusion underlying the Riemannian manifold MALA algorithm of Girolami and Calderhead (2011) is equivalent to an instance of nonlinear preconditioning. One can make a similar equivalence in the case of Riemannian manifold HMC, whereby we make a nonlinear transformation to the momentum variable used in (2), see Hoffman et al. (2019) for an explanation.
These equivalences provide motivation for further study. For instance if one can identify a $g$ such that the metric $J(g(x))^{-1}J(g(x))^{-T}$ matches that used by Girolami and Calderhead (2011) one can bypass the computationally costly operations inherent in the Riemannian manifold methods. One can also evaluate the benefits of using Riemannian schemes with arbitrary metrics by evaluating the change in the condition number under transformations which achieve those metrics.
5.3 Beyond well-conditioned distributions
The condition number as defined in (4) is restrictive in the class of models it applies to, namely distributions satisfying Assumption 2.1. Where $\Pi$ satisfies $M$-smoothness and a Poincaré inequality: for all $f\in L^{1}(\Pi)$
$\text{Var}_{\Pi}(f)\leq C_{\mathsf{PI}}\mathbb{E}_{\Pi}[\|\nabla f\|^{2}]$
with constant $C_{\mathsf{PI}}\geq 0$ Zhang et al. (2023b, footnote, page 3) define it as $\kappa:=C_{\mathsf{PI}}M$. They are motivated by its presence in the mixing time bounds they derive for the unadjusted Langevin sampler. An application of the Brascamp-Lieb inequality shows that $C_{\mathsf{PI}}=m^{-1}$ in the case that $\Pi$ also has an $m$-strongly convex potential. Chen and Gatmiry (2023) also derive mixing time bounds under a more general constraint than $m$-strong convexity. One could alternatively use the quantities involved in their constraints and therefore the mixing time bounds to redefine the condition number.
Altmeyer (2022) constructs a surrogate posterior whose potential satisfies Assumption 2.1 and coincides with the potential of the target posterior on a region in which the target concentrates. Under assumptions, they provide polynomial time mixing bounds for unadjusted Langevin Monte Carlo using the fact that the chain will stay in the aforementioned region for exponentially long with high probability. The ability to identify such behaviour allows one to quantify the conditioning of a posterior whose potential violates Assumption 2.1.
Acknowledgments and Disclosure of Funding
MH is funded by an EPSRC DTP. SL acknowledges support from EPSRC grant EP/V055380/1. The authors thank Dootika Vats, Sam Power and Giacomo Zanella for useful discussions.
Appendix A Appendix A
A.1 Proof of Proposition 1
The Hessian of the model with potential
$U(x,y)=\frac{m-M}{2}\left(\cos x+\cos y\right)+\frac{M+m}{2}\left(\frac{x^{2}}{2}+\frac{y^{2}}{2}\right)$
is in the form $\nabla^{2}U(x,y)=\text{diag}\{f(x),f(y)\}$ where $f(x):=(1/2)(M-m)\cos x+(1/2)(M+m)\in[m,M]$. As detailed in Proposition 2, the condition number is ignorant as to whether the preconditioner $L$ is symmetric or not, so we assume it is. Therefore we can perform an eigendecomposition $L=QDQ^{T}$ where $D=\text{diag}\{\lambda_{1},\lambda_{2}\}$ is the matrix of eigenvalues (not necessarily ordered) and, since we are in two dimensions, $Q$ can be represented as the two dimensional Givens matrix
$Q=\begin{pmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{pmatrix}.$
The matrix enclosed by the first operator norm in (6) has trace and determinant
$\text{Tr}(x,y)$ $:=\text{Tr}(L^{-T}\nabla^{2}U(x,y)L^{-1})=c^{2}(\lambda_{1}^{-2}f(x)+\lambda_{2}^{-2}f(y))+s^{2}(\lambda_{2}^{-2}f(x)+\lambda_{1}^{-2}f(y))$
$\text{Det}(x,y)$ $:=\text{Det}(L^{-T}\nabla^{2}U(x,y)L^{-1})=\lambda_{1}^{-2}\lambda_{2}^{-2}f(x)f(y)$
where we have abbreviated $c:=\cos\theta$, $s:=\sin\theta$ for notational simplicity. The matrix enclosed by the second operator norm in (6) has trace and determinant
$\text{Tr}^{*}(x^{*},y^{*})$ $:=\text{Tr}(L\nabla^{2}U(x^{*},y^{*})^{-1}L^{T})=c^{2}(\lambda_{1}^{2}f(x^{*})^{-1}+\lambda_{2}^{2}f(y^{*})^{-1})+s^{2}(\lambda_{2}^{-2}f(x^{*})^{-1}+\lambda_{1}^{-2}f(x^{*})^{-1})$
$\text{Det}^{*}(x^{*},y^{*})$ $:=\text{Det}(L\nabla^{2}U(x^{*},y^{*})^{-1}L^{T})=\lambda_{1}^{2}\lambda_{2}^{2}f(x^{*})^{-1}f(y^{*})^{-1}$
Using the fact that the operator norm of a positive definite matrix is simply the largest eigenvalue, we are able to lower bound
$\kappa_{L}\geq\frac{1}{2}\left(\text{Tr}(x,y)+\sqrt{\text{Tr}(x,y)^{2}-4\text{Det}(x,y)}\right)\frac{1}{2}\left(\text{Tr}^{*}(x^{*},y^{*})+\sqrt{\text{Tr}^{*}(x^{*},y^{*})^{2}-4\text{Det}^{*}(x^{*},y^{*})}\right)$
Choosing $(x,y)$ such that $f(x)=f(y)=M$ and $(x^{*},y^{*})$ such that $f(x^{*})=f(y^{*})=m$ we have
$\kappa_{L}$ $\geq\frac{1}{2}\left((\lambda_{1}^{-2}+\lambda_{2}^{-2})M+\left|\lambda_{1}^{-2}-\lambda_{2}^{-2}\right|M\right)\frac{1}{2}\left((\lambda_{1}^{2}+\lambda_{2}^{2})m^{-1}+\left|\lambda_{1}^{2}-\lambda_{2}^{2}\right|m^{-1}\right)$
$=\max\{\lambda_{1}^{-2},\lambda_{2}^{-2}\}\max\{\lambda_{1}^{2},\lambda_{2}^{2}\}\frac{M}{m}$
$=\kappa(LL^{T})\kappa$
Therefore $\kappa_{L}>\kappa$ for non-orthogonal $L$.
A.2 Proof of Proposition 2
Recall the definition of the condition number after preconditioning with a matrix $L$: $\kappa_{L}=\sup_{x}\|L^{-T}\nabla^{2}U(x)L^{-1}\|\sup_{x}\|L\nabla^{2}U(x)^{-1}L^{T}\|$. We inspect the norm inside the first supremum after computing the singular value decomposition of $L=U\Sigma V^{T}$ where $U$ and $V$ are orthogonal and $\Sigma\in\mathbb{R}^{d\times d}$ is a diagonal matrix whose diagonal elements are the singular values of $L$.
$\|L^{-T}\nabla^{2}U(x)L^{-1}\|$ $=\|U\Sigma^{-1}V^{T}\nabla^{2}U(x)V^{T}\Sigma^{-1}U\|$
$=\|\Sigma^{-1}V^{T}\nabla^{2}U(x)V^{T}\Sigma^{-1}\|$
$=\|V\Sigma^{-1}V^{T}\nabla^{2}U(x)V^{T}\Sigma^{-1}V\|$
$=\|\tilde{L}^{-T}\nabla^{2}U(x)\tilde{L}^{-1}\|$
Now we inspect the norm inside the second supremum:
$\|L\nabla^{2}U(x)^{-1}L^{T}\|$ $=\|U\Sigma V^{T}\nabla^{2}U(x)^{-1}V^{T}\Sigma U\|$
$=\|\Sigma V^{T}\nabla^{2}U(x)^{-1}V^{T}\Sigma\|$
$=\|V\Sigma V^{T}\nabla^{2}U(x)^{-1}V^{T}\Sigma V\|$
$=\|\tilde{L}\nabla^{2}U(x)^{-1}\tilde{L}^{T}\|$
A.3 Implications from Assumption 3.2
From the statement of Assumption 3.2 it is immediate that $v_{i}(x)^{T}v_{i}\geq 1-\delta$. Note that the assumption implies the following bound $\|v_{i}(x)-v_{i}\|\leq\sqrt{2}\left(1-\sqrt{1-\delta}\right)$. For $i,j\in[d]$ such that $i\neq j$, the reverse triangle inequality gives us that
$\|v_{i}(x)-v_{j}\|$ $\geq\|v_{j}-v_{i}\|-\|v_{i}-v_{i}(x)\|$
$\geq\sqrt{2}-\sqrt{2}\left(1-\sqrt{1-\delta}\right)$
and so
$\sqrt{2(1-\langle v_{i}(x),v_{j}\rangle}\geq\sqrt{2}-\sqrt{2}\left(1-\sqrt{1-\delta}\right)$
hence $v_{i}(x)^{T}v_{j}\leq\delta$ as required.
A.4 Proof of Theorem 3
Perform the eigendecomposition $\nabla^{2}U(x)=O_{x}D_{x}O_{x}^{T}$ for $O_{x}\in O(d)$ with columns $v_{i}(x)$ and $D_{x}\in\mathbb{R}^{d\times d}$ diagonal with elements $\lambda_{i}(x)$. Perform the eigendecomposition $L=V\Sigma V^{T}$ where $V$ has columns $v_{i}\in\mathbb{R}^{d}$ for $i\in[d]$ and $\Sigma:=\text{diag}\{\sigma_{1},...,\sigma_{d}\}$. Defining $\mathcal{E}_{x}:=V^{T}O_{x}-\mathbf{I}_{d}$, Assumption 3.2 guarantees that the elements of $\mathcal{E}_{x}$ are at most $\delta$ in absolute value. Inspecting the first term in the definition of $\kappa_{L}$, we have that
$\|L^{-T}\nabla^{2}U(x)L^{-1}\|$ $=\|\Sigma^{-1}(\mathcal{E}_{x}+\mathbf{I}_{d})D_{x}(\mathcal{E}_{x}+\mathbf{I}_{d})^{T}\Sigma^{-1}\|$
$\leq\|\Sigma^{-1}\mathcal{E}_{x}D_{x}\mathcal{E}_{x}^{T}\Sigma^{-1}\|+2\|\Sigma^{-1}\mathcal{E}_{x}D_{x}\Sigma^{-1}\|+\|\Sigma^{-1}D_{x}\Sigma^{-1}\|$
$\leq\|\Sigma^{-1}\mathcal{E}_{x}D_{x}\mathcal{E}_{x}^{T}\Sigma^{-1}\|+2\|\Sigma^{-1}\mathcal{E}_{x}D_{x}\Sigma^{-1}\|+(1+\epsilon)$
where the second line is due to the triangle inequality of the matrix 2-norm, the last line due to Assumption 3.1. Inspecting the norm in the second term in the above:
$\|\Sigma^{-1}\mathcal{E}_{x}D_{x}\Sigma^{-1}\|^{2}$ $=\sup_{\|v\|=1}\sum_{k=1}^{d}\left(\sum_{s=1}^{d}\frac{\lambda_{s}(x)}{\sigma_{s}\sigma_{k}}(\mathcal{E}_{x})_{ks}v_{s}\right)^{2}$
$\leq\delta^{2}\sup_{\|v\|=1}\sum_{k=1}^{d}\left(\sum_{s=1}^{d}\frac{\lambda_{s}(x)}{\sigma_{s}\sigma_{k}}v_{s}\right)^{2}$
$=\delta^{2}\sum_{k=1}^{d}\frac{1}{\sigma_{k}^{2}}\sup_{\|v\|=1}\left(\sum_{s=1}^{d}\frac{\lambda_{s}(x)}{\sigma_{s}}v_{s}\right)^{2}$
$\leq\delta^{2}(1+\epsilon)^{2}\sum_{k=1}^{d}\frac{1}{\sigma_{k}^{2}}\sup_{\|v\|=1}\left(\sum_{s=1}^{d}\sigma_{s}v_{s}\right)^{2}$
$=\delta^{2}(1+\epsilon)^{2}\sum_{k=1}^{d}\frac{1}{\sigma_{k}^{2}}\sum_{s=1}^{d}\sigma_{s}^{2}$
where the second line comes from Assumption 3.2 and the fourth line comes from Assumption 3.1. Looking at the first term now:
$\|\Sigma^{-1}\mathcal{E}_{x}D_{x}\mathcal{E}_{x}^{T}\Sigma^{-1}\|$ $=\|\Sigma^{-1}\mathcal{E}_{x}D_{x}^{\frac{1}{2}}\|^{2}$
$=\sup_{\|v\|=1}\sum_{k=1}^{d}\left(\sum_{s=1}^{d}\frac{\sqrt{\lambda_{s}(x)}}{\sigma_{k}}(\mathcal{E}_{x})_{ks}v_{s}\right)^{2}$
$\leq\delta^{2}\sum_{k=1}^{d}\frac{1}{\sigma_{k}^{2}}\sup_{\|v\|=1}\left(\sum_{s=1}^{d}\sqrt{\lambda_{s}(x)}v_{s}\right)^{2}$
$\leq\delta^{2}(1+\epsilon)\sum_{k=1}^{d}\frac{1}{\sigma_{k}^{2}}\sup_{\|v\|=1}\left(\sum_{s=1}^{d}\sigma_{s}v_{s}\right)^{2}$
$=\delta^{2}(1+\epsilon)\sum_{k=1}^{d}\frac{1}{\sigma_{k}^{2}}\sum_{s=1}^{d}\sigma_{s}^{2}$
where the third line comes from Assumption 3.2 and the fourth line comes from Assumption 3.1. Putting the terms together yields
$\|L^{-T}\nabla^{2}U(x)L^{-1}\|\leq(1+\epsilon)\left(1+\delta\sqrt{\sum_{i=1}^{d}\sigma_{i}^{2}\sum_{i=1}^{d}\sigma_{i}^{-2}}\right)^{2}$
Now we follow the same procedure for $\|L\nabla^{2}U(x)L^{T}\|$:
$\|L\nabla^{2}U(x)L^{T}\|$ $\leq\|\Sigma\mathcal{E}_{x}D_{x}^{-\frac{1}{2}}\|^{2}+2\|\Sigma\mathcal{E}_{x}D_{x}^{-1}\Sigma\|+\|\Sigma D_{x}^{-1}\Sigma\|$
$\leq\|\Sigma\mathcal{E}_{x}D_{x}^{-\frac{1}{2}}\|^{2}+2\|\Sigma\mathcal{E}_{x}D_{x}^{-1}\Sigma\|+(1+\epsilon)$
starting with the second term:
$\|\Sigma\mathcal{E}_{x}D_{x}^{-1}\Sigma\|^{2}$ $=\sup_{\|v\|=1}\sum_{k=1}^{d}\left(\sum_{s=1}^{d}\frac{\sigma_{s}\sigma_{k}}{\lambda_{s}(x)}(\mathcal{E}_{x})_{ks}v_{s}\right)^{2}$
$\leq\delta^{2}\sum_{k=1}^{d}\sigma_{k}^{2}\sup_{\|v\|=1}\left(\sum_{s=1}^{d}\frac{\sigma_{s}}{\lambda_{s}(x)}v_{s}\right)^{2}$
$\leq\delta^{2}(1+\epsilon)^{2}\sum_{k=1}^{d}\sigma_{k}^{2}\sup_{\|v\|=1}\left(\sum_{s=1}^{d}\frac{1}{\sigma_{s}}v_{s}\right)^{2}$
$\leq\delta^{2}(1+\epsilon)^{2}\sum_{k=1}^{d}\sigma_{k}^{2}\sum_{s=1}^{d}\frac{1}{\sigma_{s}^{2}}$
and the first term:
$\|\Sigma\mathcal{E}_{x}D_{x}^{-\frac{1}{2}}\|^{2}$ $=\sup_{\|v\|=1}\sum_{k=1}^{d}\left(\sum_{s=1}^{d}\frac{\sigma_{k}}{\sqrt{\lambda_{s}(x)}}(\mathcal{E}_{x})_{ks}v_{s}\right)^{2}$
$\leq\delta^{2}\sum_{k=1}^{d}\sigma_{k}^{2}\sup_{\|v\|=1}\left(\sum_{s=1}^{d}\frac{1}{\sqrt{\lambda_{s}(x)}}v_{s}\right)^{2}$
$\leq\delta^{2}(1+\epsilon)\sum_{k=1}^{d}\sigma_{k}^{2}\sup_{\|v\|=1}\left(\sum_{s=1}^{d}\frac{1}{\sigma_{s}}v_{s}\right)^{2}$
$=\delta^{2}(1+\epsilon)\sum_{k=1}^{d}\sigma_{k}^{2}\sum_{s=1}^{d}\frac{1}{\sigma_{s}^{2}}$
from which follows
$\kappa_{L}\leq(1+\epsilon)^{2}\left(1+\delta\sqrt{\sum_{i=1}^{d}\sigma_{i}^{2}\sum_{i=1}^{d}\sigma_{i}^{-2}}\right)^{4}$
A.5 Tight $\delta$ dependence in Theorem 3
Take the first norm of (4): $\|L^{-T}\nabla^{2}U(x)L^{-1}\|=\|Q_{\pi}GD_{\pi}^{1/2}G^{T}Q_{\pi}^{T}Q_{\pi}D_{\pi}^{-1}Q_{\pi}^{T}Q_{\pi}GD_{\pi}^{1/2}G^{T}Q_{\pi}^{T}\|=\|D_{\pi}^{1/2}G^{T}D_{\pi}^{-1}GD_{\pi}^{1/2}\|$. Similarly the second norm is $\|D_{\pi}^{-1/2}G^{T}D_{\pi}GD_{\pi}^{-1/2}\|$. Therefore $\kappa_{L}=\lambda_{1}(M)\lambda_{2}(M)^{-1}$ where $M=D_{\pi}^{1/2}G^{T}D_{\pi}^{-1}GD_{\pi}^{1/2}$. Since $G$ is a perturbation of angle $\arccos(1-\delta)$ it has the form
$G=\begin{pmatrix}1-\delta&-\sqrt{\delta(2-\delta)}\\ \sqrt{\delta(2-\delta)}&1-\delta\end{pmatrix}$
and so
$M=\begin{pmatrix}(1-\delta)^{2}+\delta(2-\delta)\frac{\lambda_{1}}{\lambda_{2}}&\left(\sqrt{\frac{\lambda_{1}}{\lambda_{2}}}-\sqrt{\frac{\lambda_{2}}{\lambda_{1}}}\right)(1-\delta)\sqrt{\delta(2-\delta)}\\ \left(\sqrt{\frac{\lambda_{1}}{\lambda_{2}}}-\sqrt{\frac{\lambda_{2}}{\lambda_{1}}}\right)(1-\delta)\sqrt{\delta(2-\delta)}&(1-\delta)^{2}+\delta(2-\delta)\frac{\lambda_{2}}{\lambda_{1}}\end{pmatrix}$
where $\lambda_{1}$ and $\lambda_{2}$ are the diagonal elements of $D_{\pi}$. We have that $\text{tr}(M)=2(1-\delta)^{2}+\delta(2-\delta)(\lambda_{1}\lambda_{2}^{-1}+\lambda_{1}^{-1}\lambda_{2})$ and $\det(M)=1$. That the determinant is one means that $\kappa_{L}=\lambda_{1}(M)^{2}$. Using the trace-determinant formulation of the eigenvalues of a $2\times 2$ matrix we have that
$\lambda_{1}(M)=\frac{1}{2}\left(2(1-\delta)^{2}+\delta(2-\delta)\left(\frac{\lambda_{1}}{\lambda_{2}}+\frac{\lambda_{2}}{\lambda_{1}}\right)+\sqrt{\left(2(1-\delta)^{2}+\delta(2-\delta)\left(\frac{\lambda_{1}}{\lambda_{2}}+\frac{\lambda_{2}}{\lambda_{1}}\right)\right)^{2}-4}\right)$
so, unless $\lambda_{1}=\lambda_{2}$ and $\Sigma_{\pi}$ is a multiple of the identity, $\kappa_{L}=O(\delta^{4})$.
A.6 Assumption 3.3 implies Assumption 3.1
Weyl’s inequality implies that
$\frac{\lambda_{i}(\nabla^{2}U(x))}{\sigma_{i}^{2}}\leq\frac{\lambda_{i}(LL^{T})+\lambda_{1}(\nabla^{2}U(x)-LL^{T})}{\sigma_{i}^{2}}\leq 1+\frac{\|\nabla^{2}U(x)-LL^{T}\|}{\sigma_{i}^{2}}$
and so $\|\nabla^{2}U(x)-LL^{T}\|\leq\sigma_{d}^{2}\epsilon$ implies Assumption 3.1 with the same $\epsilon$.
A.7 Proof of Theorem 4
Based on the intuition gained from Proposition 2 we can assume that $L$ is symmetric, and so its left and right singular vectors are simply its eigenvectors. Using (Yu et al., 2015, Corollary 1) with $\hat{\Sigma}=\nabla^{2}U(x)$ and $\Sigma=LL^{T}$ we have that $\|v_{i}(x)-v_{i}\|\leq 2^{\frac{3}{2}}\gamma^{-1}\|\nabla^{2}U(x)-LL^{T}\|$. Rearranging, the Assumption 3.3 gives $\langle v_{i}(x),v_{i}\rangle\geq 1-4\gamma^{-2}\sigma_{d}^{-4}\epsilon^{2}$. From A.6, Assumption 3.3 gives us Assumption 3.1 with the same $\epsilon$, and hence we can apply Theorem 3 with $\delta=1-(1-2\gamma^{-1}\sigma_{d}^{-2}\epsilon)^{2}$.
A.8 Proof of Theorem 5
Using Proposition 2 we assume that $L$ is symmetric. For the first supremum in the definition of $\kappa_{L}$ note that $\|\nabla^{2}U(x)-L^{2}\|=\|L^{T}(L^{-T}\nabla^{2}U(x)L^{-1}-\mathbf{I}_{d})L\|$. Using the fact that $\sigma_{i}(BA)\leq\|B\|\sigma_{i}(A)$ and $\sigma_{i}(AC)\leq\sigma_{i}(A)\|C\|$ for matrices $A,B,C$ of appropriate sizes and all $i\in[d]$ (Tao, 2012, exercise 1.3.24) we have
$\|L^{T}(L^{-T}\nabla^{2}U(x)L^{-1}-\mathbf{I}_{d})L\|$ $\geq\frac{\sigma_{1}(L^{T}(L^{-T}\nabla^{2}U(x)L^{-1}-\mathbf{I}_{d})LL^{-1})}{\sigma_{1}(L^{-1})}$
$=\sigma_{d}(L)\|L^{T}(L^{-T}\nabla^{2}U(x)L^{-1}-\mathbf{I}_{d})\|$
$\geq\sigma_{d}(L)\frac{\sigma_{1}(L^{-T}L^{T}(L^{-T}\nabla^{2}U(x)L^{-1}-\mathbf{I}_{d}))}{\sigma_{1}(L^{-T})}$
$=\sigma_{d}(L)^{2}\|L^{-T}\nabla^{2}U(x)L^{-1}-\mathbf{I}_{d}\|$
Therefore we can bound $\|L^{-T}\nabla^{2}U(x)L^{-1}-\mathbf{I}_{d}\|\leq\epsilon$ using Assumption 3.3. Using the reverse triangle inequality $\|L^{-T}\nabla^{2}U(x)L^{-1}-\mathbf{I}_{d}\|\geq|\|L^{-T}\nabla^{2}U(x)L^{-1}\|-1|$ we get $\|L^{-T}\nabla^{2}U(x)L^{-1}\|\leq 1+\epsilon$.
For the second supremum in the definition of $\kappa_{L}$ we use the same technique as the first supremum, first noting that $\|\nabla^{2}U(x)^{-1}-L^{-2}\|\leq\|\nabla^{2}U(x)^{-1}\|\|L^{-2}\|\|\nabla^{2}U(x)-L^{2}\|\leq m^{-1}\epsilon$. Employing the technique from before:
$\|\nabla^{2}U(x)^{-1}-L^{-2}\|$ $=\|L^{-1}(L\nabla^{2}U(x)^{-1}L^{T}-\mathbf{I}_{d})L^{-T}\|$
$\geq\sigma_{d}(L^{-1})^{2}\|L\nabla^{2}U(x)^{-1}L^{T}-\mathbf{I}_{d}\|$
and hence $\|L\nabla^{2}U(x)^{-1}L^{T}-\mathbf{I}_{d}\|\leq\sigma_{1}(L)^{2}m^{-1}\epsilon$. Using the reverse triangle inequality again gives $\|L\nabla^{2}U(x)^{-1}L^{T}\|\leq 1+\sigma_{1}(L)^{2}m^{-1}\epsilon$.
A.9 Proof of Proposition 6
Applying the non-rectangular form of Ostrowski’s theorem (Higham and Cheng, 1998a, Theorem 3.2) gives for any $x\in\mathbb{R}^{d}$
$\lambda_{d}(B^{T}B)\lambda_{n-d+1}(\Lambda(x))\leq\lambda_{1}(B^{T}\Lambda(x)B)\leq\lambda_{1}(B^{T}B)\lambda_{1}(\Lambda(x)),$
and similarly
$\lambda_{d}(B^{T}B)\lambda_{n}(\Lambda(x))\leq\lambda_{d}(B^{T}\Lambda(x)B)\leq\lambda_{1}(B^{T}B)\lambda_{d}(\Lambda(x))$
Since $\kappa:=\sup_{x\in\mathbb{R}^{d}}\lambda_{1}(B^{T}\Lambda(x)B)/\inf_{x\in\mathbb{R}^{d}}\lambda_{d}(B^{T}\Lambda(x)B)$ then applying the upper/lower bound to $\lambda_{1}(B^{T}\Lambda(x)B)$ and the lower/upper bound to $\lambda_{d}(B^{T}\Lambda(x)B)$ point-wise gives the upper/lower bound on $\kappa$ as desired.
A.10 Proof of Proposition 8
Setting $\tilde{X}^{T}=(X^{T}X)^{-1/2}X^{T}$ and applying Proposition 6 gives the result, noting that $\tilde{X}^{T}\tilde{X}=\mathbf{I}_{d}$.
A.11 Proof of Proposition 10
First note that the preconditioned Hessian can be written
$L^{-T}\nabla^{2}U(x)L^{-1}=(X^{T}\Lambda(x^{*})X)^{1/2}X^{T}\Lambda(x^{*})^{1/2}\Lambda(x^{*})^{-1/2}\Lambda(x)\Lambda(x^{*})^{-1/2}\Lambda(x^{*})^{1/2}X(X^{T}\Lambda(x^{*})X)^{1/2}$
Setting $\tilde{X}^{T}:=(X^{T}\Lambda(x^{*})X)^{1/2}X^{T}\Lambda(x^{*})^{1/2}$ and then applying the upper bound of Proposition 6 to the matrix $\tilde{X}^{T}\Lambda(x^{*})^{-1/2}\Lambda(x)\Lambda(x^{*})^{-1/2}\tilde{X}$ gives the first inequality. The second follows from applying the same bound again to $\Lambda(x^{*})^{-1/2}\Lambda(x)\Lambda(x^{*})^{-1/2}$ and noting that
$\kappa(\Lambda(x^{*}))\leq\frac{\sup_{x\in\mathbb{R}^{d}}\lambda_{1}(\Lambda(x))}{\inf_{x\in\mathbb{R}^{d}}\lambda_{d}(\Lambda(x))}$
A.12 Proof of Theorem 11
If we take $\sigma^{2}=\xi/(Md)$, then (Andrieu et al., 2022, Theorem 1) implies that the spectral gap $\gamma_{\kappa}$ of the RWM algorithm on a target with a $m$-strongly convex, $M$-smooth potential is bounded as follows:
$C\xi\exp(-2\xi)\frac{1}{\kappa}\frac{1}{d}\leq\gamma_{\kappa}\leq\frac{\xi}{2}\frac{1}{d}$
We will modify the proof of (Andrieu et al., 2022, Lemma 47) so that the upper bound on $\gamma_{\kappa}$ subsequently depends on $\kappa$. The spectral gap of the RWM algorithm on a target $\pi$ with kernel $P$ is defined as $\gamma_{k}:=\inf_{f\in L_{0}^{2}(\pi)}(\mathcal{E}(P,f)/\text{Var}_{\pi}(f))$ where $\mathcal{E}$ is the Dirichlet form associated with $\pi$. Define $g(x):=\langle v_{\text{max}},x-\mathbb{E}_{\pi}[X]\rangle$ where $v_{\text{max}}\in\mathbb{R}^{d}$ is the eigenvector associated with the greatest eigenvalue of $E_{\pi}\left(\nabla^{2}U(X)\right)^{-1}$. The Cramér-Rao inequality gives that
$\text{Var}_{\pi}(g(X))$ $\geq v_{\text{max}}^{T}\mathbb{E}_{\pi}\left[\nabla^{2}U(X)\right]^{-1}v_{\text{max}}$
$=\lambda_{1}\left(\mathbb{E}_{\pi}\left[\nabla^{2}U(X)\right]^{-1}\right)\|v_{\text{max}}\|^{2}$
$=\lambda_{d}\left(\mathbb{E}_{\pi}\left[\nabla^{2}U(X)\right]\right)^{-1}\|v_{\text{max}}\|^{2}$
The second equality comes from the fact that $\mathbb{E}_{\pi}\left[\nabla^{2}U(X)\right]^{-1}$ is positive definite, since it is the inverse of the expectation of a matrix that is itself positive definite.
Say $v\in\mathbb{R}^{d}$ is the eigenvector associated with the smallest eigenvalue $\lambda_{d}(y)$ of $\nabla^{2}U(y)$ for a given $y\in\mathbb{R}^{d}$. Then
$\lambda_{d}\left(\mathbb{E}_{\pi}\left[\nabla^{2}U(X)\right]\right)$ $=\inf_{\|v\|=1}v^{T}\mathbb{E}_{\pi}\left[\nabla^{2}U(X)\right]v$
$\leq v^{T}\mathbb{E}_{\pi}\left[\nabla^{2}U(X)\right]v$
$=\mathbb{E}_{\pi}\left[v^{T}\left(\nabla^{2}U(X)-\nabla^{2}U(y)+\nabla^{2}U(y)\right)v\right]$
$\leq\sup_{x\in\mathbb{R}^{d}}\|\nabla^{2}U(x)-\nabla^{2}U(y)\|+\lambda_{d}(y)$
$\leq m(1+2\epsilon)$
where in the final line we use Assumption 3.6, and the fact that Assumption 3.6 implies
$\frac{\lambda_{d}(y)}{\lambda_{d}(x)}\leq 1+\epsilon$
for all $x,y\in\mathbb{R}^{d}$ (see A.6) and hence $\lambda_{d}(y)\leq(1+\epsilon)\lambda_{d}(x)\leq(1+\epsilon)m$. Therefore $\text{Var}_{\pi}(g(X))\geq m^{-1}(1+2\epsilon)^{-1}\|v_{\text{max}}\|^{2}$. Upper bounding the Dirichlet form in the same way as (Andrieu et al., 2022, Lemma 47) gives $\mathcal{E}(P,g)\leq(1/2)\sigma^{2}\|v_{\text{max}}\|^{2}$, and so
$\gamma_{k}=\inf_{f\in L_{0}^{2}(\pi)}\frac{\mathcal{E}(P,f)}{\text{Var}_{\pi}(f)}\leq\frac{\mathcal{E}(P,g)}{\text{Var}_{\pi}(g)}\leq\frac{\frac{1}{2}\sigma^{2}\|v_{\text{max}}\|^{2}}{m^{-1}(1+2\epsilon)^{-1}\|v_{\text{max}}\|^{2}}=\frac{1}{2}\xi\kappa^{-1}d^{-1}(1+2\epsilon)$
A.13 Proof of Corollary 12
The lower bound for the spectral gap post-preconditioning is $\gamma_{\kappa_{L}}\geq C\xi\exp(-2\xi)\kappa_{L}^{-1}d^{-1}$ due to (Andrieu et al., 2022, Theorem 1). The target satisfies Assumption 3.3 so we can use Theorem 5 to modify the bound: $\gamma_{\kappa_{L}}\geq C\xi\exp(-2\xi)(1+\epsilon)^{-1}\left(1+m^{-1}\sigma_{1}(L)^{2}\epsilon\right)^{-1}d^{-1}$. So, applying the upper bound found in Theorem 11 to the spectral gap before preconditioning, we see that a condition number $\kappa$ such that
$\frac{1}{2}\xi\kappa^{-1}d^{-1}(1+2\epsilon^{\prime})\leq C\xi\exp(-2\xi)(1+\epsilon)^{-1}\left(1+\frac{\sigma_{1}(L)^{2}}{m}\epsilon\right)^{-1}d^{-1}$
guarantees that $\gamma_{\kappa_{L}}\geq\gamma_{\kappa}$ and so increases the spectral gap.
A.14 Proof of Proposition 15
The preconditioned O-U process has generator $A_{L}=\text{tr}\left(L^{-1}L^{-T}\nabla^{2}\right)+\left\langle-L^{-1}L^{-T}\Sigma_{\pi}^{-1}x,\nabla\right\rangle$ for $x\in\mathbb{R}^{d}$, where $\nabla$ is the grad operator and
$\nabla^{2}:=\sum_{i=1}^{d}\frac{\partial^{2}}{\partial x_{i}^{2}}$
The spectral gap of a generator of a stochastic process with generator $A$ is defined as the smallest distance from its spectrum (excluding 0) to 0.
Denoting spectrum of $-L^{-1}L^{-T}\Sigma_{\pi}^{-1}$ by $\lambda_{1},...,\lambda_{d}$ Metafune et al. (2002, Theorem 3.1) has that the spectrum of the generator $A_{L}$ is $\{\lambda=\sum_{i=1}^{d}n_{i}\lambda_{i}:n_{i}\in\mathbb{N}\}$. Therefore the spectral gap is
$\min\left\{\left|\sum_{i=1}^{d}n_{i}\lambda_{i}\right|:n_{i}\in\mathbb{N}\backslash\{0\}\text{ for }i\in[d]\right\}=\min_{i\in[d]}\left|\lambda_{i}\right|$
Say $\left|\lambda_{i}\right|>1$ for some $i$ without loss of generality. Then the constraint $\left|\det(-L^{-1}L^{-T}\Sigma_{\pi}^{-1})\right|=1$ implies that $\left|\lambda_{j}\right|<1$ for some $j\neq i$, and hence the resulting spectral gap is strictly less than one. In the case that $\left|\lambda_{i}\right|=1$ for all $i$, the spectral gap is exactly one, and therefore optimal. This is achieved when $L=\Sigma_{\pi}^{-1/2}$ for any notion of the matrix square root.
A.15 Proof of Proposition 16
We assume WLOG that $U(x^{*})=0$. Taylor’s theorem with integral remainder has that
[…]
(since $U(x^{*})=0$, $\nabla U(x^{*})=0$) from which we can deduce
$\frac{1}{2}(x-x^{*})^{T}\Delta_{-}(x-x^{*})\leq U(x)\leq\frac{1}{2}(x-x^{*})^{T}\Delta_{+}(x-x^{*})$
and hence
$\exp\left(-\frac{1}{2}(x-x^{*})^{T}\Delta_{+}(x-x^{*})\right)\leq\exp(-U(x))\leq\exp\left(-\frac{1}{2}(x-x^{*})^{T}\Delta_{-}(x-x^{*})\right)$
with
$\frac{Z_{\Delta_{+}}}{Z}\frac{1}{Z_{\Delta_{+}}}\exp\left(-\frac{1}{2}(x-x^{*})^{T}\Delta_{+}(x-x^{*})\right)\leq\frac{1}{Z}\exp(-U(x))\leq\frac{Z_{\Delta_{-}}}{Z}\frac{1}{Z_{\Delta_{-}}}\exp\left(-\frac{1}{2}(x-x^{*})^{T}\Delta_{-}(x-x^{*})\right)$
where $Z_{A}:=\sqrt{(2\pi)^{d}\det A^{-1}}$. For an arbitrary $v\in\mathbb{R}^{d}$ we have that
[…]
$\leq\frac{Z_{\Delta_{-}}}{Z}\mathbb{E}_{\mathcal{N}(x^{*},(\Delta_{-})^{-1})}[(v^{T}(X-\mu_{\pi}))^{2}]$
$\leq\frac{Z_{\Delta_{-}}}{Z}\left(\mathbb{E}_{\mathcal{N}(x^{*},(\Delta_{-})^{-1})}[(v^{T}(X-x^{*}))^{2}]-(v^{T}(x^{*}-\mu_{\pi}))^{2}\right)$
$\leq\frac{Z_{\Delta_{-}}}{Z_{\Delta_{+}}}\left(v^{T}(\Delta_{-})^{-1}v-v^{T}(x^{*}-\mu_{\pi})(x^{*}-\mu_{\pi})^{T}v\right)$
where the last inequality follows from the fact that $Z_{\Delta_{+}}\leq Z\leq Z_{\Delta_{-}}$. We can construct a similar lower bound to give
$c((\Delta_{+})^{-1}-(x^{*}-\mu_{\pi})(x^{*}-\mu_{\pi})^{T})\leq\Sigma_{\pi}\leq c^{-1}((\Delta_{-})^{-1}-(x^{*}-\mu_{\pi})(x^{*}-\mu_{\pi})^{T})$
defining $c:=(Z_{\Delta_{+}}/Z_{\Delta_{-}})=\sqrt{\det\Delta_{-}\det\Delta_{+}^{-1}}\leq 1$. This gives $P_{-}\leq\Sigma_{\pi}^{-1}\leq P_{+}$ where
$P_{+}$ $:=c^{-1}\left((\Delta_{+})^{-1}-(x^{*}-\mu_{\pi})(x^{*}-\mu_{\pi})^{T}\right)^{-1}$
$P^{-}$ $:=c\left((\Delta_{-})^{-1}-(x^{*}-\mu_{\pi})(x^{*}-\mu_{\pi})^{T}\right)^{-1}$
and hence
$P_{+}$ $=c^{-1}\left(\Delta_{+}+\left(1-(x^{*}-\mu_{\pi})^{T}\Delta_{+}(x^{*}-\mu_{\pi})\right)^{-1}\Delta_{+}(x^{*}-\mu_{\pi})(x^{*}-\mu_{\pi})^{T}\Delta_{+}\right)$
$P_{-}$ $=c\left(\Delta_{-}+\left(1-(x^{*}-\mu_{\pi})^{T}\Delta_{-}(x^{*}-\mu_{\pi})\right)^{-1}\Delta_{-}(x^{*}-\mu_{\pi})(x^{*}-\mu_{\pi})^{T}\Delta_{-}\right)$
using the Woodbury identity. The fact that $(x^{*}-\mu_{\pi})^{T}\Delta_{\pm}(x^{*}-\mu_{\pi})=\text{Tr}(D_{\pm})$ gives the result.
We have that $\|\nabla^{2}U(x)-\Sigma_{\pi}^{-1}\|:=\sup_{v}\left|v^{T}\nabla^{2}U(x)v-v^{T}\Sigma_{\pi}^{-1}v\right|$. Say the quantity inside the absolute value is positive. Then $v^{T}\nabla^{2}U(x)v-v^{T}\Sigma_{\pi}^{-1}v\leq v^{T}\nabla^{2}U(x)v-v^{T}P_{-}v$. Now say the quantity is negative, giving us $v^{T}\Sigma_{\pi}^{-1}v-v^{T}\nabla^{2}U(x)v\leq v^{T}P_{+}v-v^{T}\nabla^{2}U(x)v$. In sum this gives
$\|\nabla^{2}U(x)-\Sigma_{\pi}^{-1}\|$ $\leq\sup_{v:\|v\|=1}\max\left\{v^{T}\nabla^{2}U(x)v-v^{T}P_{-}v,v^{T}P_{+}v-v^{T}\nabla^{2}U(x)v\right\}$
$\leq\sup_{v:\|v\|=1}\max\left\{v^{T}\Delta_{+}v-v^{T}P_{-}v,v^{T}P_{+}v-v^{T}\Delta_{-}v\right\}$
$\leq\max\left\{\|\Delta_{+}-P_{-}\|,\|P_{+}-\Delta_{-}\|\right\}$
A.16 Proof of Proposition 17
From proposition 16 we have that $P_{-}\leq\Sigma_{\pi}^{-1}\leq P_{+}$where
$P_{+}$ $=c^{-1}\left(A+\epsilon\mathbf{I}_{d}+\left(1-(x^{*}-\mu_{\pi})^{T}(A+\epsilon\mathbf{I}_{d})(x^{*}-\mu_{\pi})\right)^{-1}(A+\epsilon\mathbf{I}_{d})(x^{*}-\mu_{\pi})(x^{*}-\mu_{\pi})^{T}(A+\epsilon\mathbf{I}_{d})\right)$
$P_{-}$ $=c\left(A-\epsilon\mathbf{I}_{d}+\left(1-(x^{*}-\mu_{\pi})^{T}(A-\epsilon\mathbf{I}_{d})(x^{*}-\mu_{\pi})\right)^{-1}(A-\epsilon\mathbf{I}_{d})(x^{*}-\mu_{\pi})(x^{*}-\mu_{\pi})^{T}(A-\epsilon\mathbf{I}_{d})\right)$
since $\Delta_{-}=A-\epsilon\mathbf{I}_{d}$ and $\Delta_{+}=A+\epsilon\mathbf{I}_{d}$. The bounds stated at the end of the proposition give
$\|\nabla^{2}U(x)-\Sigma_{\pi}^{-1}\|$ $\leq\max\left\{\|\Delta_{+}-P_{-}\|,\|P_{+}-\Delta_{-}\|\right\}$
$=\max\left\{\|(1-c)A+(1+c)\epsilon\mathbf{I}_{d}-c\tilde{P}_{-}\|,\|(c^{-1}-1)A+(c^{-1}+1)\epsilon\mathbf{I}_{d}-c^{-1}\tilde{P}_{+}\|\right\}$
$\leq(c^{-1}-1)\|A\|+(c^{-1}+1)\epsilon+\max\left\{c\|\tilde{P}_{-}\|,c^{-1}\|\tilde{P}_{+}\|\right\}$
where in the final line we use the triangle inequality.